### v1

# SenseShift6D: Multimodal RGB-D Benchmarking for Robust 6D Pose Estimation across Environment and Sensor Variations

Abstract
Recent advances on 6D object-pose estimation has achieved high performance on representative benchmarks such as LM-O, YCB-V, and T-Less. However, these datasets were captured under fixed illumination and camera settings, leaving the impact of real-world variations in illumination, exposure, gain or depth-sensor mode – and the potential of test-time sensor control to mitigate such variations – largely unexplored. To bridge this gap, we introduce SenseShift6D, the first RGB-D dataset that physically sweeps 13 RGB exposures, 9 RGB gains, auto-exposure, 4 depth-capture modes, and 5 illumination levels. For three common household objects (spray, pringles, and tincase), we acquire 101.9k RGB and 10k depth images, which can provide 1,380 unique sensor-lighting permutations per object pose. Experiments with state-of-the-art models on our dataset show that applying sensor control during test-time induces greater performance improvement over digital data augmentation, achieving performance comparable to or better than costly increases in real-world training data quantity and diversity. Adapting either RGB or depth sensors individually is effective, while jointly adapting multimodal RGB-D configurations yields even greater improvements. SenseShift6D extends the 6D-pose evaluation paradigm from data-centered to sensor-aware robustness, laying a foundation for adaptive, self-tuning perception systems capable of operating robustly in uncertain real-world environments.
Our dataset is available at: huggingface.co/datasets/Yegyu/SenseShift6D
Associated scripts can be found at: github.com/yegyu-han/SenseShift6D
1 Introduction
Figure 1: RGB sample images under auto exposure and under all combinations of exposure and gain settings for brightness 5%, 25%, 75% and 100%. Rows indicate gain levels and columns indicate exposure levels.
Figure 2: Depth sample images under four depth capture modes: default, high accuracy, high density, and medium density.
Estimating the 6D pose of everyday objects is a critical computer vision task, enabling fine-grained interactions in mixed-reality headsets, autonomous-vehicle manipulation, and embodied AI agents navigating cluttered environments  [5, 28, 24, 26, 16, 21]. The field has made substantial progress, with state-of-the-art methods achieving high accuracy on established benchmarks [7, 3, 10, 15, 11].
However, these impressive results are achieved under strictly controlled acquisition settings. Real-world deployments, in contrast, face variations: ambient illumination changes throughout the day, camera exposure and gain settings are adjusted by on-board auto-control or user preference, and commodity depth sensors can also switch capture modes on purpose. These environmental and sensor dynamics introduce complex artifacts, including nonlinear color shifts and structured depth noise, which are challenging to replicate accurately with synthetic augmentations or compensate for during post-processing [1]. Consequently, existing benchmarks – collected with fixed illumination, exposure, gain, and depth configurations – cannot adequately represent real-world scenarios. Therefore, the research community lacks a systematic benchmark for quantifying how 6D-pose estimators respond to per-frame sensor variation or for evaluating emerging test-time sensor-control strategies [1, 2] designed to mitigate these shifts.
Recent attempts to tackle this robustness gap have explored adaptive sensor control, which dynamically adjusts sensor parameters at test time to maximize accuracy without modifying the underlying model. Studies in image classification have demonstrated that test-time sensor control improves model robustness more effectively than scaling up the model size or augmenting training data [1, 2]. However, these efforts have been limited to unimodal RGB tasks. Importantly, no prior work has examined multimodal (RGB and depth) sensor control or provided a benchmark that physical sweeps multimodal sensor parameters with ground-truth 6D pose. Consequently, the effectiveness of adaptive multimodal sensing and its interplay with traditional data-centric training remain an open question.
To bridge this gap, we introduce SenseShift6D, a novel RGB-D benchmark for 6D pose estimation that physically sweeps key sensor and lighting parameters. Using an Intel RealSense RGB-D camera in a controlled darkroom setup, we systematically vary 13 exposure levels, 9 gain settings, 4 depth-capture modes, and 5 illumination conditions across three representative household objects: a textured spray can, a cylindrical Pringles tube, and a glossy tin case. This protocol produces a dataset of 101.9k RGB and 10k depth frames, which can provide 1,380 unique sensor-lighting permutations per object pose (as illustrated in Figures 1 and 2), each annotated via 6D ground-truth via ArUco calibration and manual refinement (with $\sim$3 mm errors). For controlled study, experiments in this paper are conducted on the four subsets of the entire dataset: Train-Def, recorded under default camera and illumination settings, and Train-Var, covering a comprehensive multimodal parameter grid (8 exposure, 5 gain, auto-exposure, 4 depth-capture settings with 5 lighting conditions). Correspondingly, we construct two test splits (Test-Def and Test-Var) without overlap with the training configurations. As all variations are physically realized, the captured images preserve authentic sensor artifacts, including nonlinear color shifts, photon-shot noise, and structured depth errors, that synthetic augmentations cannot replicate. Thus, SenseShift6D establishes the first realistic testbed for sensor-aware 6D pose estimation and adaptive sensing.
This paper offers three main contributions:
  • Sensor-aware multimodal benchmarking: We present the first 6D object-pose benchmark that systematically and physically varies multimodal sensor parameters across frames, enabling fine-grained analysis of sensor-dependent performance degradation.
  • Robustness through test-time sensor control: We demonstrate that adjusting sensor parameters during inference under environmental variations improves 6D pose estimation accuracy by an average of 11.4% compared to standard auto-exposure with conventional digital augmentation. This indicates that sensor-level tuning can better account for the real-world variability faced during deployment. Moreover, sensor control proves to be a data-efficient approach to improving robustness, achieving comparable generalization with significantly less training data.
  • Potential of multimodal sensor adaptation: While adapting either RGB or depth sensors individually already yields noticeable improvements, jointly optimizing both RGB and depth parameters at test-time provides even greater gains. This emphasizes the value of studying multimodal sensor interactions and highlights new directions for advancing the field of 6D pose estimation.
By providing a structured testbed for sensor-aware object 6D pose estimation, SenseShift6D lays the groundwork for more robust, adaptive perception systems in dynamic real-world environments.
2 Related Work
Table 1: Comparison of 6D pose estimation datasets in terms of sensor and lighting variation.
Dataset Environmental and sensor variations Objects RGB frames Depth frames
Illumination Exposure Gain Depth
LM (ACCV ’12) [7] $\times$ $\times$ $\times$ $\times$ 15 18.2k 18.2k
LM-O (ECCV ’14) [3] $\times$ $\times$ $\times$ $\times$ 8 1.2k 1.2k
T-LESS (WACV ’17) [10] $\times$ $\times$ $\times$ $\times$ 30 49k 49k
YCB-V (RSS ’18) [27] $\times$ $\times$ $\times$ $\times$ 21 133k 133k
TUD-L (ECCV ’18) [11] 8 locations $\times$ $\times$ $\times$ 3 62k 62k
HB (ICCVW ’19) [15] 2 locations $\times$ $\times$ $\times$ 33 34.8k 34.8k
HOPE (IROS ’22) [23] 5 locations $\times$ $\times$ $\times$ 28 238 238
IPD (CVPR ’24) [14] 3 locations 4 levels $\times$ $\times$ 20 30k 30k
SenseShift6D (Ours) 5 levels 13 levels 9 levels 4 modes 3 101.9k 10k
2.1 6D Pose Estimation Benchmark
Classic datasets. LineMOD (LM) [7] and its occlusion subset (LM-O) [3] pioneered RGB-D pose estimation benchmarks in controlled tabletop setups. Subsequent benchmarks such as T-LESS [10], which provided a large collection of texture-less industrial objects, and YCB-V [27], which offered cluttered scenes with temporal sequences, expanded the scope of evaluation. While these benchmarks provide robust evaluation regarding variations in objects and their poses, all were captured under fixed illumination and camera settings.
Initial exploration of environmental and sensor variation. TUD-L [11], HomebrewedDB (HB) [15], and HOPE [23] introduced variations in illumination conditions but retained fixed camera settings with auto-exposure enabled. The recent Industrial Parts Dataset (IPD) [14] was the first benchmark to vary camera parameters, capturing four exposure levels under multiple lighting conditions. However, IPD focuses on texture-less industrial parts and does not explore variations in RGB gain or depth-sensor modes, thus limiting its suitability for adaptive-sensing research.
Remaining gap and our contribution. As summarized in Table 1, no existing benchmark offers orthogonal sweeps of RGB exposure, RGB gain, and multiple depth-capture modes together with dense 6D ground truth. Our SenseShift6D dataset directly addresses this gap by systematically varying exposure, gain, and depth-capture modes under diverse illumination conditions, enabling fine-grained analysis of 6D pose estimation robustness to realistic hardware-level changes.
2.2 Adaptive Sensor Control Benchmark
Recent studies have explored test-time adaptation of camera sensor parameters to improve downstream vision performance. ImageNet-ES [1] and ImageNet-ES-Diverse [2] captured real-world scenes using a physical camera, systematically varying aperture, shutter speed, and ISO across multiple lighting conditions for image classification tasks. These benchmarks demonstrated that physical variation cannot be replicated by synthetic augmentations and dynamically selecting optimal camera parameters at test time can significantly increase the performance of lightweight models, enabling accuracy competitive with larger foundation models without additional training.
However, existing adaptive sensing research remains limited to single-modal RGB classification. To date, no prior work has explored: (1) multimodal sensor control, where depth-sensor modes must be jointly optimized with RGB settings; or (2) the impact of adaptive sensor control on object-centric 6D pose estimation tasks, which inherently rely on both photometric and geometric cues. Our SenseShift6D addresses this gap by providing the first realistic testbed for evaluating multimodal, per-frame sensor-control policies alongside conventional data-centric training baselines.
3 SenseShift6D
Table 2: SenseShift6D Configurations for Data split.
Data split Brightness (%) Sensor configuration Exposure ($\mu$s) Gain Depth-capture mode Captured images
Train-Def 50 Auto (1 RGB option, 1 depth option) Auto Auto Default RGB: 385, Depth: 385 1 RGB-D scene/pose
Train-Var 5, 25, 50, 75, 100 Auto + Manual (41 RGB options, 1 depth options) Auto, 2, 4, 19, 78, 312, 1250, 5000, 10000 Auto, 0, 32, 64, 96, 128 Default RGB: 80.0k, Depth: 1.9k 205 RGB-D scenes/pose
Test-Def 50 Auto (1 RGB option, 1 depth option) Auto Auto Default RGB: 128, Depth: 128 1 RGB-D scenes/pose
Test-Var 5, 25, 50, 75, 100 Auto + Manual (21 RGB options, 4 Depth options) Auto, 9, 39, 156, 625, 2500 Auto, 16, 48, 80, 112 Default, High Accuracy, High Density, Medium Density RGB: 13k, Depth: 2.5k 420 RGB-D scenes/pose
Figure 3: Three household objects collected in SenseShift6D.
We introduce SenseShift6D, the first RGB-D benchmark that physically sweeps camera and lighting parameters, enabling rigorous evaluation of 6D object pose estimators under realistic environmental and sensor dynamics. Using an Intel RealSense D455 mounted in a darkroom, we systematically capture three representative household objects – a textured spray can, a cylindrical Pringles tube, and a glossy tin case (Figure 3) – under auto-exposure, 13 RGB-exposure times, 9 RGB-gain settings, 4 depth-capture presets, and 5 illumination levels (details in Table 2). Our dataset consists of 101.9k RGB and 10k depth frames, which can yield 1,380 unique sensor-lighting permutations per object pose. Accurate ground truth is obtained via a rigidly mounted ChArUco board and careful hand-eye calibration.
To facilitate controlled benchmarking, we organize the dataset into four clearly defined splits:
  • Train-Def: Contains only the default sensor and lighting configuration (1 RGB-D scene/pose)
  • Train-Var: Expands Train-Def, covering (8 exposures $\times$ 5 gains + 1 auto-exposure) $\times$ 1 depth modes $\times$ 5 brightness levels (205 RGB-D scenes/pose).
  • Test-Def: Includes the default sensor and lighting configuration (1 RGB-D scenes/pose)
  • Test-Var: Expands Test-Def, featuring (5 exposures $\times$ 4 gains + 1 auto-exposure) configurations, spanning the same depth-mode and illumination range (420 RGB-D scenes/pose).
As all variations are physically captured rather than digitally simulated, the resulting images inherently retain authentic artifacts, such as non-linear color shifts, photon-shot noise, and depth-mode-specific errors, that synthetic augmentation methods cannot reproduce [1]. SenseShift6D therefore offers the first realistic testbed for analyzing environment- and sensor-aware robustness, enabling side-by-side evaluation of per-frame adaptive-sensing policies against traditional data-centric training strategies.
3.1 Environmental and Sensor Variations
Ambient-illumination sweep. Real deployments frequently encounter lighting changes, such as those found in warehouse aisles, outdoor loading bays, or dim domestic environments. To reproduce these diverse conditions in a controlled and repeatable manner, we built a darkroom enclosed with black curtains to block external light completely. Within this setup, we installed three Philips Hue White & Color Ambiance bulbs, providing precise and programmable brightness control. Each bulb was calibrated to a consistent color temperature of 6535 K, with brightness systematically varied across five levels: 5%, 25%, 50%, 75% and 100%. This controlled setting ensures uniform illumination conditions, enabling accurate comparisons of RGB gain and exposure settings under identical photon flux.
RGB-sensor parameters. The Intel RealSense D455 camera permits manual control of two key RGB sensor parameters: exposure, determining the duration of sensor exposure to incoming light, and gain, which amplifies the captured signal strength.
  • Exposure grid. The D455 supports exposure settings ranging from 1 $\mu$s to 10,000 $\mu$s. Empirical tests indicated that exposure times below 1,000 $\mu$s yield image quality comparable to the camera’s auto-exposure setting in well-lit environments. Given that this threshold is relatively low within the configurable range, we chose 13 logarithmically spaced exposure values to maximize coverage across practical operating conditions.
  • Gain grid. The RGB gain of the D455 can be set from 0 to 128. Through visual inspection, we found that linear steps produce more diverse characteristics than exponential spacing. Therefore, we selected 9 linearly spaced gain settings spanning the full range.
Depth capture modes. The Intel RealSense D455 generates depth maps via stereo matching, estimating depth by aligning left and right camera images. This process relies on multiple internal parameters whose individual effects are complex and challenging to fine-tune directly. To manage this complexity, we leverage four predefined depth presets provided by Intel, each optimized for different usage scenarios [12]:
  • Default: Offers visually clean depth maps with reduced noise and well-defined edges, suitable for general-purpose applications.
  • High Accuracy: Provides highly reliable depth estimates by enforcing stricter confidence thresholds, albeit at the expense of a lower fill rate.
  • High Density: Increases the fill factor, providing more complete depth maps, particularly effective for detecting more surfaces in low-texture areas.
  • Medium Density: Offers a balanced trade-off between fill factor and accuracy, aiming to balance depth map completeness and precision.
3.2 Dataset Collection
Figure 4: Testbed for data collection.
Table 3: GT annotation accuracy (mm).
$\mu_{\delta}$ $\sigma_{\delta}$ $\mu_{|\delta|}$ $med_{|\delta|}$
RealSense -2.15 4.55 3.75 2.89
To capture each scene systematically, we acquired images under all sensor configurations associated with a given data split, while maintaining fixed positions for the object and camera. Scenes were assigned exclusively to specific splits, ensuring no overlap between training and test sets. To capture diverse object poses, we used a motorized turntable to vary object orientation and adjusted the camera’s position and tilt angle. Each object was rigidly mounted onto a ChArUco board, which itself was firmly placed on the turntable, which provides a consistent and precise pose reference throughout data capture.
For automatic collection, the turntable was remotely controlled, and deliberate delays were introduced after changing lighting brightness or sensor parameters, allowing sufficient stabilization time. We employed one smart IR remote-control hub to send commands to the turntable and another hub to control three smart bulbs. Both hubs received commands from a cloud-based server, which coordinated the entire data-capture capture sequence through an integrated control script (see Figure 4).
3.3 Data Processing
Validation. We manually reviewed the images for each parameter configuration. If any scene appeared inconsistent due to lighting or sensor parameter change delays, we excluded that scene across all configurations to maintain dataset consistency.
Stabilizing auto-exposure frames. The default setting to capture RGB images is the auto-exposure mode. For the D455 camera, we observed that the auto-exposure setting depends on the gain value set in the previous frame, leading to frame-to-frame brightness jitter. To suppress this effect, we captured successive auto-exposure shots of each scene (5 shots for Train-Def/Var and 4 shots for Test-Def/Var), each preceded by a different gain initialization, and average them to obtain a stable reference image.
Ground-Truth annotation. For each pose, we used the 50%-brightness, averaged auto-exposure RGB frame and went through the annotation process below, resulting in 2$\sim$3 mm errors.
  • Object masks: Initially, rough bounding boxes were generated using Grounding DINO [18], guided by shape-aware textual prompts tailored to each object. Boxes were manually reviewed and corrected if needed. The refined boxes were passed to FastSAM [29] to generate precise instance-level masks.
  • Object poses: An initial camera-to-object transformation was provided by detecting the ChArUco markerboard rigidly attached to the turntable. The object poses were then refined precisely in 3D using the annotation utility [6] in the BOP toolkit [9] with depth maps captured in High Density mode. Finally, all refined poses were re-projected onto the RGB images and visually inspected to confirm alignment accuracy.
  • Accuracy of GT annotation: Following the protocol in [10, 15], we calculate the difference $\delta=d_{c}-d_{r}$, where $d_{c}$ denotes the captured depth map and $d_{r}$ denotes the rendered depth map using the GT pose. Differences over 5 cm are treated as outliers and omitted. Table 4 reports the mean $\mu_{\delta}$, standard deviation $\sigma_{\delta}$, and the mean and median of the absolute differences $\mu_{|\delta|}$ and $med_{|\delta|}$. High-accuracy depth sensing is employed for reliable evaluation of GT accuracy analysis.
4 Experiments
To assess how sensor control improves efficiency in 6D pose estimation relative to conventional training and inference approaches, we conducted experiments on several state-of-the-art models.
4.1 Experiment Setup
Models. We chose three 6D object pose estimation models: ZebraPose [22], GDRNPP [19], and HiPose [17]. ZebraPose is an RGB-based method that leverages features extracted from a ResNet-34 backbone. GDRNPP improves upon its predecessor, GDR-Net  [25], by adopting a ConvNeXt  [20] backbone and modifying training details such as learning rate scheduling and weight decay. While GDRNPP includes a depth-based pose refinement module, its inference primarily relies on RGB inputs. HiPose is an RGB-D model that uses ConvNeXt as the RGB backbone and RandLA-Net  [13] to extract point cloud features from depth. These RGB and depth features are fused through a bidirectional interaction module for pose estimation. In our experiments, we excluded the refinement stage for ZebraPose and GDRNPP, while evaluating the full pipeline of HiPose.
Train setup. For training, we first generated 10k photorealistic (PBR) synthetic images per object using BlenderProc [4]. Each model was trained independently on a per-object basis, following one of the three schemes: (1) PBR + Train-Def (default physical images only), (2) PBR + Train-Def with data augmentation (including Gaussian blur, linear contrast adjustment, brightness enhancement and others – see Appendix A for details), and (3) PBR + Train-Var (full sensor and lighting variation). ZebraPose and GDRNPP were trained using an NVIDIA RTX 3090 GPU, while HiPose was trained on an NVIDIA A100 GPU. During training, we utilized depth images captured under the default-capture mode.
Metric. We used the standard Average Distance (ADD) metric [8] to evaluate model performance. ADD measures the average point-wise distance between model points transformed by the predicted pose and the ground-truth pose:
$\text{ADD}=\frac{1}{m}\sum_{\mathbf{x}\in\mathcal{M}}\left\|(\mathbf{R}\mathbf{x}+\mathbf{T})-(\tilde{\mathbf{R}}\mathbf{x}+\tilde{\mathbf{T}})\right\|,$ (1)
where $\mathcal{M}$ denotes the set of 3D model points, $\mathbf{x}\in\mathcal{M}$ represents an individual point on the model, $m$ is the total number of points in the model. $\mathbf{R}$ and $\mathbf{T}$ are the ground-truth rotation matrix and translation vector, while $\tilde{\mathbf{R}}$ and $\tilde{\mathbf{T}}$ are their predicted counterparts.
4.2 Analysis on RGB Sensor Control
Table 4 evaluates the impact of illumination and RGB sensor variations on 6D pose estimation. All reported results utilize the ADD metric, using AR@5, which corresponds to recall at a threshold of 5% of the object diameter. We compare three sensor control settings: (1) AE (auto-exposure) represents the conventional baseline with automatic sensor settings; (2) Rand (random exposure and gain) indicates average performance across randomly selected sensor configurations; (3) Oracle represents the best achievable performance through optimal sensor control – measuring the probability that at least one sensor-controlled configuration enables correct pose prediction for a given scene.
Table 4: Performance of various sensor control methods under different training and test settings of SenseShift6D.
ZebraPose [22] GDRNPP [19] HiPose [17]
Test-Def Test-Var Test-Def Test-Var Test-Def Test-Var
Object Train AE AE Rand Oracle AE AE Rand Oracle AE AE Rand Oracle
Spray PBR + Train-Def 100.0 96.09 52.26 99.51 100.0 96.59 52.02 99.02 100.0 98.73 60.00 100.0
PBR + Train-Def w/ Aug 100.0 98.04 68.46 100.0 100.0 98.54 71.88 100.0 100.0 100.0 89.61 100.0
PBR + Train-Var 97.56 97.07 78.80 97.56 97.56 97.56 85.93 98.54 100.0 100.0 100.0 100.0
Pringles PBR + Train-Def 85.71 75.71 38.87 94.75 69.05 60.95 35.64 89.05 100.0 96.00 58.74 100.0
PBR + Train-Def w/ Aug 73.80 66.18 40.92 89.99 80.95 79.05 49.17 89.52 97.62 97.81 72.81 100.0
PBR + Train-Var 90.47 89.04 68.35 95.23 90.48 90.004 72.22 97.14 97.62 97.62 86.86 98.10
Tincase PBR + Train-Def 95.55 84.43 44.15 96.00 82.22 75.11 43.29 90.67 73.33 60.36 40.27 77.78
PBR + Train-Def w/ Aug 95.55 88.88 53.37 99.10 73.33 69.78 48.85 94.67 75.56 72.44 51.09 76.00
PBR + Train-Var 97.77 96.88 73.55 99.55 95.56 95.11 77.05 97.78 75.56 73.87 73.09 75.11
Overall PBR + Train-Def 93.75 85.41 45.09 96.75 72.60 77.53 43.65 92.91 91.11 85.03 53.00 92.59
PBR + Train-Def w/ Aug 89.78 84.37 54.25 96.36 79.20 82.45 56.63 94.73 91.06 90.08 71.17 92.00
PBR + Train-Var 95.26 94.33 73.56 97.44 94.53 94.23 78.40 97.82 91.06 90.50 86.65 91.07
In Table 4, AE consistently exhibits lower performance on Test-Var compared to Test-Def (captured at the same 50% brightness level as Train-Def), regardless of the model or training strategy. This clearly illustrates the necessity of evaluating model robustness under realistic illumination shifts. Applying digital augmentation during training generally improves AE performance on Test-Var, yet adversely affects certain cases (e.g., ZebraPose for Pringles and GDRNPP for Tincase), suggesting that augmentation alone may not reliably enhance robustness. In contrast, training models on physically collected data (Train-Var, which is 3.5 times larger than Train-Def) consistently yields greater performance gains compared to digital augmentation, highlighting the inherent limitations of synthetic simulation in accurately replicating real-world environmental variations. Given that physically collecting large-scale, diverse datasets is often challenging in practice, this performance gap underscores the need for an efficient test-time adaptation methods to effectively address sensor and environmental shifts.
Model-friendly sensor control vs. Traditional auto-exposure. Within Test-Var, AE significantly outperforms Rand, showing the critical role of intentional sensor control in managing environmental variability. Furthermore, Oracle consistently surpasses AE across all models and training strategies, verifying that the auto-exposure mechanism – typically optimized for human visual perception – is not necessarily optimal for model performance. Thus, substantial potential exists for model-specific sensor adaptation at test time, without modifying model architectures or training schemes. Notably, Oracle achieves over 13.4% average performance improvement compared to AE for RGB-based models (ZebraPose and GDRNPP).
Test-time sensor control vs. Traditional generalization approaches. Oracle without augmentation (i.e., the first training scheme) consistently surpasses AE with augmentation, and even outperforms AE trained on Train-Var in most scenarios. In the worst case (GDRNPP), Oracle trained on Train-Def remains comparable to AE trained on Train-Var, showing less than a 2% accuracy gap. These superior results, relative to traditional generalization strategies, such as augmentation or manual scaling training data, demonstrate the substantial potential of test-time sensor control for 6D pose estimation in dynamic real-world environments. Consequently, mitigating covariate shifts during inference through sensor optimization may be more effective, or complementary, compared to exclusively focusing on model generalization to all possible covariate variations, as currently pursued by foundational-model approaches.
4.3 Analysis on Multimodal RGB-D Sensor Control
To comprehensively analyze the impact of multimodal sensor control under environmental variations, we train HiPose (an RGB-D model) following the first training scheme (PBR + Train-Def, captured at 50% brightness level) and evaluate it on the Test-Var split. To specifically assess high-precision pose estimation, we report AUC@[0:0.1], which measures the area under the ADD recall curve within 10% of the object diameter. We consider five conditions: (1) Baseline utilizes the default settings for both RGB and depth sensors; (2) Depth-Only applies Oracle selection for the depth sensor, while maintaining auto-exposure settings for RGB; (3) RGB-Only applies Oracle selection for the RGB sensor, while using the Default mode for the depth sensor; (4) Oracle-Fixed selects the single best-performing fixed RGB-D configuration and applies it across all test scenes; (5) Oracle-Dynamic applies Oracle selection to both RGB and depth sensors for each scene.
Table 5: AUC performance of various sensor control methods for HiPose, trained on PBR + Train-Def and tested on Test-Var.
Object Baseline RGB: Auto Depth: Default Depth-Only RGB: Auto Depth: Oracle RGB-Only RGB: Oracle Depth: Default Oracle-Fixed Best Fixed Param. Oracle-Dynamic RGB: Oracle Depth: Oracle
Spray 86.80 88.04(+1.24) 88.87(+2.07) 86.80 89.92(+3.12)
Pringles 73.65 76.10(+2.45) 78.92(+5.27) 74.68(+1.03) 81.42(+7.77)
Tincase 56.07 58.63(+2.56) 63.60(+7.53) 56.10(+0.03) 66.92 (+10.85)
Overall 72.17 74.25(+2.08) 77.13(+4.95) 72.53(+0.53) 79.42(+7.24)
Table 5 shows that all four sensor-control methods outperform Baseline across every scenario. Among these methods, Oracle-Fixed yields the smallest improvement, offering only marginal gains over Baseline despite controlling multimodal sensor parameters. This result highlights the impracticality of relying on a single, fixed sensor configuration to handle diverse environmental variations; instead, sensor parameters must be adapted dynamically on a per-scene basis.
Complementary role of RGB and Depth. Interestingly, Depth-Only achieves a notable AUC improvement of 2.08% compared to Baseline, demonstrating consistent benefits across all objects. Even though the model was trained exclusively with the Default depth-capture mode – suggesting that employing alternative modes at inference might introduce covariate shifts – allowing flexible depth-mode selection still substantially enhances pose estimation performance. RGB-Only further improves performance relative to Depth-Only, particularly for symmetric objects (Pringles and Tincase). This suggests that RGB surface information may provide crucial cues for accurately estimating the poses of symmetric objects.
Multimodal joint control vs. Unimodal control. Oracle-Dynamic achieves the best overall performance, improving AUC by 7.24% over Baseline and by 2.95% over the second-best RGB-Only method. Notably, Oracle-Dynamic exhibits the most significant gain for Tincase (10.85% above Baseline), the most challenging object in our dataset. These results indicate that RGB and depth modalities provide complementary information, and that jointly adapting both sensors creates synergy exceeding that achievable by independently adapting individual modalities. Thus, our findings highlight the importance of deeper investigations into multimodal sensor adaptation strategies and understanding which combinations of RGB-D inputs are most beneficial for model performance.
4.4 Qualitative Analysis
Analysis on RGB sensor control. Figure 5 illustrates two representative test scenes for GDRNPP, involving the Tincase and Pringles objects, respectively. In both cases, AE results in incorrect pose predictions (via GDRNPP), while Oracle (RGB-only) successfully estimates accurate object poses. These scenes were captured at the lowest brightness level, representing the most challenging environmental condition. These examples highlight that optimal RGB sensor parameters, especially gain values in these cases, vary depending on both object type and object pose, underscoring the necessity of per-scene test-time sensor control. Furthermore, the RGB images selected by Oracle and AE appear visually indistinguishable from a human perspective. This indicates that the optimal direction for model-friendly sensor control is not always intuitively aligned with human visual perception, emphasizing the importance of sensor control strategies specifically tailored to model performance rather than subjective human assessment.
Analysis on depth sensor control. Figure 6 illustrates two representative test scenes evaluated with HiPose, involving the Tincase and Pringles objects, respectively. Both scenes were captured at the lowest brightness level, representing the most challenging environmental conditions. While the RGB sensor parameters were fixed, we compared pose estimation performance across different depth-capture modes. In both cases, the Default depth mode resulted in incorrect pose predictions, whereas Oracle (Depth-only) successfully produced accurate object poses. These examples clearly demonstrate that different depth-capture modes yield varying depth-map qualities, affecting pose estimation outcomes. Furthermore, the optimal depth-capture mode depends on the specific object type and scene configuration, underscoring the importance of adaptive, scene-dependent depth-sensor control strategies.
Figure 5: RGB visualization of pose estimation results under auto exposure and sensor control settings. The red line represents the ground truth pose, and the green line indicates the predicted pose.
Figure 6: Visualization of pose estimation results showing that depth modes other than the default yield better performance under a fixed RGB configuration (E2500, G16). Red lines denote ground truth poses, while green lines represent predicted poses.
5 Conclusion
We introduced SenseShift6D, the first sensor-aware, multimodal benchmark for 6D object-pose estimation that physically sweeps RGB exposure, RGB gain, depth-capture mode, and ambient illumination. Extensive experiments show that state-of-the-art methods suffer substantial accuracy drops under these realistic sensor shifts. Crucially, we demonstrated that adaptive sensor control at test time can recover a large portion of this lost performance – without any model retraining – and that joint RGB + depth control outperforms unimodal strategies. These findings highlight sensor control as a promising and lightweight avenue for deploying 6D-pose systems in dynamic, real-world conditions.
Limitations and Future Work. SenseShift6D currently focuses on single-object tabletop scenes and varies only brightness intensity, omitting complexities such as directional lighting, shadows, and occlusions. Future extensions of the dataset will include: (1) multi-object scenarios to explore occlusion and clutter; (2) directional and shadow-rich lighting conditions; and (3) a real-time adaptive sensor-control framework that dynamically adjusts RGB and depth parameters. We believe SenseShift6D provides a strong foundation for developing robust, sensor-adaptive 6D pose estimation systems suitable for deployment in diverse and dynamic real-world environments.
Appendix
Appendix A RGB Augmentation Settings
To improve generalization under varying environmental conditions, we adopted an RGB augmentation pipeline based on that used in GDRNPP. The augmentations were implemented using the imgaug library and applied only to the RGB input during training with a probability of 0.8. An overview of these augmentations is illustrated in Figure 7.
The augmentations listed in Table 6 were applied in random order, with each operation activated independently with a fixed probability.
Table 6: Augmentation operations and parameters.
Augmentation Type Activation Probability Parameters / Range
Gaussian Blur 0.4 $\sigma\in[0.0,3.0]$
Sharpness Enhancement 0.3 factor $\in[0.0,50.0]$
Contrast Enhancement 0.3 factor $\in[0.2,50.0]$
Brightness Enhancement 0.5 factor $\in[0.1,6.0]$
Color Enhancement 0.3 factor $\in[0.0,20.0]$
Additive Intensity 0.5 value $\in[-25,25]$, per-channel = 0.3
Invert Pixels 0.3 probability = 0.2, per-channel
Multiply Intensity 0.5 factor $\in[0.6,1.4]$, per-channel = 0.5
Additive Gaussian Noise 0.1 scale = 10, per-channel
Linear Contrast 0.5 factor $\in[0.5,2.2]$, per-channel = 0.3
Grayscale Conversion 0.5 $\alpha\in[0.0,1.0]$
Figure 7: Examples of RGB augmentations applied during training. The top-left image is the original, and the remaining images are randomly augmented samples generated using the settings summarized in Table 6.
Appendix B Licenses
We use open-source components under various licenses for evaluation and data generation. In particular, some scripts for photorealistic rendering are adapted from BlenderProc (GPL-3.0), and evaluation pipelines integrate components from GDRNPP (Apache-2.0), ZebraPose (MIT), and HiPose (MIT). All third-party code is used in compliance with their respective licenses, with attribution provided as required.
Our dataset, SenseShift6D, is distributed under the Creative Commons Attribution 4.0 International (CC BY 4.0) license.
Appendix C Experimental Details
C.1 Compute Resources
All models were trained on a high-performance computing cluster equipped with NVIDIA GPUs and Intel Xeon CPUs. Specifically, ZebraPose and GDRNPP were trained on nodes with RTX 3090 GPUs (24GB VRAM) and Intel Xeon Silver 4210R CPUs, while HiPose was trained on a node with A100 GPUs (80GB VRAM) and an Intel Xeon Gold 6530 CPU. Each training node had access to 512GB of RAM.
C.2 Model Training
All models were trained with a batch size of 32. ZebraPose and HiPose were both trained for a total of 38,000 iterations using the Adam optimizer, with learning rates of 0.0002 and 0.0001, respectively. GDRNPP was trained for 120 epochs using the Ranger optimizer with a learning rate of 8e-4 and weight decay of 0.01. A flat-and-anneal learning rate schedule with cosine annealing was applied, starting at 72% of the total training epochs. A linear warmup was used for the first 1,000 iterations with a warmup factor of 0.001. All learning rate schedules and optimizer settings follow the original implementations of each method.
Table 7: Configurations for various data split.
Data split Brightness (%) Sensor configuration Exposure ($\mu$s) Gain Depth-capture mode
Train-B5Def 5 Auto (1 RGB option, 1 depth option) Auto Auto Default
Train-BADef 5, 25, 50, 75, 100 Auto (1 RGB option, 1 depth option) Auto Auto Default
Train-B50Var 50 Auto + Manual (41 RGB options, 1 depth options) Auto, 2, 4, 19, 78, 312, 1250, 5000, 10000 Auto, 0, 32, 64, 96, 128 Default
Test-B5Def 5 Auto (1 RGB option, 1 depth option) Auto Auto Default, High Accuracy, High Density, Medium Density
Test-B5Var 5 Auto + Manual (21 RGB options, 4 Depth options) Auto, 9, 39, 156, 625, 2500 Auto, 16, 48, 80, 112 Default, High Accuracy, High Density, Medium Density
Appendix D Additional Evaluation Results
Table 8 presents the performance of the Tincase in Table 4, categorized by brightness level. This suggests a noticeable difference in performance at brightness 5 compared to other brightness levels. Therefore, we evaluated SenseShift6D dataset on various data splits, beyond those described in Table 2. Each split is detailed in Table 7.
Table 8: Performance of sensor control across brightness levels for Tincase.
ZebraPose GDRNPP HiPose
Brightness Test-Var PBR+ Train-Def PBR+ Train-Def w/ Aug PBR+ Train-Var PBR+ Train-Def PBR+ Train-Def w/ Aug PBR+ Train-Var PBR+ Train-Def PBR+ Train-Def w/ Aug PBR+ Train-Var
B5 AE 33.33 64.44 95.55 37.78 40.00 93.33 10.67 64.89 71.56
Rand 9.67 15.56 36.89 13.56 19.44 42.55 10.78 18.67 69.11
Oracle 80.00 97.78 97.78 88.89 91.11 100.00 75.56 71.11 71.11
B25 AE 97.78 93.33 95.56 84.44 73.33 95.56 71.11 73.33 73.33
Rand 46.11 51.11 70.00 43.89 46.22 75.11 38.67 44.56 73.33
Oracle 100.00 97.78 100.00 91.11 95.56 97.78 77.78 73.33 73.33
B50 AE 95.56 95.56 97.78 82.22 75.56 95.56 72.44 75.11 73.78
Rand 56.11 64.67 87.00 55.33 58.00 88.34 53.00 60.89 73.56
Oracle 100.00 100.00 100.00 91.11 97.78 97.78 77.78 77.78 75.56
B75 AE 97.78 93.33 97.78 84.44 80.00 95.56 73.78 74.22 76.44
Rand 54.22 69.44 89.78 50.22 62.78 92.11 46.78 69.44 75.33
Oracle 100.00 100.00 100.00 91.11 95.56 97.78 80.00 80.00 77.78
B100 AE 97.78 97.78 97.78 86.67 80.00 95.56 73.78 74.67 74.22
Rand 54.67 66.11 84.11 53.45 57.78 87.11 52.11 61.89 74.11
Oracle 100.00 100.00 100.00 91.11 93.33 95.56 77.78 77.78 77.78
Overall AE 84.43 88.88 96.88 75.11 69.78 95.11 60.36 72.44 73.87
Rand 44.15 53.37 73.55 42.29 48.85 77.05 40.27 51.09 73.09
Oracle 96.00 99.10 99.55 90.67 94.67 97.78 77.78 76.00 75.11
D.1 Trained without Environmental Shift
We conducted an experiment under the low-light condition of brightness 5, which yielded the lowest performance in brightness-wise evaluation. To eliminate environmental shifts, we set both training and testing to brightness 5, replacing the default training condition (brightness 50). As shown in Table 9, the oracle with sensor control achieves performance comparable to or even better than training and testing conducted entirely within the same brightness 5 domain, suggesting that adaptive sensor control can mitigate domain gaps without the need for retraining in the target environment.
Table 9: Comparison of models trained without environmental shift under brightness 5.
Model Train-B5Def Train-Def
Test-B5Def AE Test-B5Def AE Test-B5Var Rand Test-B5Var Oracle
ZebraPose 26.66 33.33 9.66 80.00
GDRNPP 86.67 37.78 13.56 88.89
HiPose 35.44 10.67 10.78 75.56
D.2 Trained on Diverse Sensor and Environmental Settings
In contrast to Train-Var, we partially incorporated sensor control and environmental variability into the training process. Table 10 presents AUC performance at [0:0.1], indicating that including sensor-controlled images at a fixed brightness level (Train-B50Var) can lead to improved generalization across diverse environmental conditions, when compared to using auto exposure images from multiple brightness levels (Train-BADef). This indicates that sensor control introduces rich variations which act as implicit regularization against domain shifts. Notably, even without additional training, appropriate sensor configuration at test time contributes significantly to model performance, as shown by the oracle performance under the Train-Def setting.
Table 10: AUC performance under various brightness and sensor configurations for Tincase.
ZebraPose GDRNPP HiPose
Test-Def Test-Var Test-Def Test-Var Test-Def Test-Var
Train AE AE Rand Oracle AE AE Rand Oracle AE AE Rand Oracle
Train-Def 76.80 68.14 38.20 84.62 67.34 60.93 31.51 77.50 61.58 55.90 34.78 64.22
Train-BADef 78.73 70.09 37.31 83.96 81.54 80.38 48.83 89.04 61.82 55.36 34.32 63.73
Train-B50Var 78.91 77.00 56.16 85.58 81.17 79.69 62.03 87.48 61.82 61.54 60.47 63.00
Train-Var 80.44 79.54 61.91 86.60 79.32 78.64 64.20 85.46 61.73 61.61 60.86 63.13
Appendix E Additional Qualitative Results on Depth Variations
Figure 8 visualizes how the object region appears in the depth map under each depth capture mode for all three objects. In addition, Figure 9 shows qualitative examples for Tincase, using HiPose trained on Train-Def. Given a fixed RGB image with poor visibility, the figure compares predictions and corresponding masked depth maps between the default mode and oracle-selected modes, where the default mode fails across all examples while the oracle modes often produce better results including some successful predictions.
Figure 8: Visualization of masked object depth maps across different capture modes.
Figure 9: Comparison of predictions under default and oracle capture modes with a fixed RGB image. (Left) Visualized object pose on RGB images: ground truth pose in red, predicted pose in green. (Right) Masked depth maps corresponding to each capture mode.

### v2

@@
 Abstract
-Recent advances on 6D object-pose estimation has achieved high performance on representative benchmarks such as LM-O, YCB-V, and T-Less.
-However, these datasets were captured under fixed illumination and camera settings, leaving the impact of real-world variations in illumination, exposure, gain or depth-sensor mode – and the potential of test-time sensor control to mitigate such variations – largely unexplored.
+Recent advances on 6D object-pose estimation have achieved high performance on representative benchmarks such as LM-O, YCB-V, and T-Less.
+However, these datasets were captured under fixed illumination and camera settings, leaving the impact of real-world variations in illumination, exposure, gain or depth-sensor mode—and the potential of test-time sensor control to mitigate such variations—largely unexplored.
 To bridge this gap, we introduce SenseShift6D, the first RGB-D dataset that physically sweeps 13 RGB exposures, 9 RGB gains, auto-exposure, 4 depth-capture modes, and 5 illumination levels.
-For three common household objects (spray, pringles, and tincase), we acquire 101.9k RGB and 10k depth images, which can provide 1,380 unique sensor-lighting permutations per object pose.
-Experiments with state-of-the-art models on our dataset show that applying sensor control during test-time induces greater performance improvement over digital data augmentation, achieving performance comparable to or better than costly increases in real-world training data quantity and diversity.
-Adapting either RGB or depth sensors individually is effective, while jointly adapting multimodal RGB-D configurations yields even greater improvements.
-SenseShift6D extends the 6D-pose evaluation paradigm from data-centered to sensor-aware robustness, laying a foundation for adaptive, self-tuning perception systems capable of operating robustly in uncertain real-world environments.
-Our dataset is available at: huggingface.co/datasets/Yegyu/SenseShift6D
-Associated scripts can be found at: github.com/yegyu-han/SenseShift6D
+For five common household objects (spray, pringles, tincase, sandwich, and mouse), we acquire 166.4k RGB and 16.7k depth images, which can provide 1,380 unique sensor-lighting permutations per object pose.
+Experiments with state-of-the-art models on our dataset demonstrate that applying multimodal sensor control at test time yields substantial performance gains, achieving a 19.5 pp improvement on pretrained generalizable models.
+It also enhances robustness precisely where those models tend to fail.
+Moreover, even instance-level pose estimators, where train and test set share identical object and background, performance still varies under environmental and sensor change, demonstrating that test-time sensor control remains effective compared to costly expansions in the quantity and diversity of real-world training data, without any additional training.
+SenseShift6D extends the object pose evaluation paradigm from data-centered to sensor-aware robustness, laying a foundation for adaptive, self-tuning perception systems capable of operating robustly in uncertain real-world environments.
 1 Introduction
+Estimating the 6D pose of everyday objects is a critical computer vision task, enabling fine-grained interactions in mixed-reality headsets, autonomous-vehicle manipulation, and embodied AI agents navigating cluttered environments  (Doughty & Ghugre, 2022; Zhang et al., 2023; Trojak et al., 2025; Wu et al., 2019; Ke et al., 2020; Ponimatkin et al., 2025).
+The field has made substantial progress, with state-of-the-art methods achieving high accuracy on established benchmarks (Hinterstoisser et al., 2012a; Brachmann et al., 2014; Hodan et al., 2017; Kaskman et al., 2019; Hodan et al., 2018b).
+However, these impressive results are achieved under strictly controlled acquisition settings.
+Real-world deployments, in contrast, face substantial variations: ambient illumination changes throughout the day; camera exposure and gain are adjusted by on-board auto-control or user preference; and commodity depth sensors switch capture modes on purpose.
+These environmental and sensor dynamics introduce complex artifacts—nonlinear color shifts and structured depth noise—that are difficult to reproduce via synthetic augmentations or compensate for during post-processing (Baek et al., 2024).
+Consequently, existing benchmarks—mostly collected with fixed illumination, exposure, gain, and depth configurations—cannot adequately represent real-world scenarios.
+Therefore, the research community lacks a systematic benchmark to quantify how 6D-pose estimators respond to per-frame sensor variation or to evaluate emerging test-time sensor-control strategies (Baek et al., 2024; 2025) designed to mitigate these shifts.
+Recent attempts to tackle this robustness gap have explored adaptive sensor control, which dynamically adjusts sensor parameters at test time to maximize accuracy without modifying the underlying model.
+Studies in image classification have demonstrated that test-time sensor control improves model robustness more effectively than scaling up the model size or augmenting training data (Baek et al., 2024; 2025).
+However, these efforts have been limited to unimodal RGB tasks.
+Importantly, no prior work has examined multimodal (RGB and depth) sensor control or provided a benchmark that physically sweeps multimodal sensor parameters with ground-truth 6D pose.
+Consequently, the effectiveness of adaptive multimodal sensing—and its interplay with traditional data-centric training—remains an open question.
 Figure 1: RGB sample images under auto exposure and under all combinations of exposure and gain settings for brightness 5%, 25%, 75% and 100%.
@@
 Figure 2: Depth sample images under four depth capture modes: default, high accuracy, high density, and medium density.
-Estimating the 6D pose of everyday objects is a critical computer vision task, enabling fine-grained interactions in mixed-reality headsets, autonomous-vehicle manipulation, and embodied AI agents navigating cluttered environments  [5, 28, 24, 26, 16, 21].
-The field has made substantial progress, with state-of-the-art methods achieving high accuracy on established benchmarks [7, 3, 10, 15, 11].
-However, these impressive results are achieved under strictly controlled acquisition settings.
-Real-world deployments, in contrast, face variations: ambient illumination changes throughout the day, camera exposure and gain settings are adjusted by on-board auto-control or user preference, and commodity depth sensors can also switch capture modes on purpose.
-These environmental and sensor dynamics introduce complex artifacts, including nonlinear color shifts and structured depth noise, which are challenging to replicate accurately with synthetic augmentations or compensate for during post-processing [1].
-Consequently, existing benchmarks – collected with fixed illumination, exposure, gain, and depth configurations – cannot adequately represent real-world scenarios.
-Therefore, the research community lacks a systematic benchmark for quantifying how 6D-pose estimators respond to per-frame sensor variation or for evaluating emerging test-time sensor-control strategies [1, 2] designed to mitigate these shifts.
-Recent attempts to tackle this robustness gap have explored adaptive sensor control, which dynamically adjusts sensor parameters at test time to maximize accuracy without modifying the underlying model.
-Studies in image classification have demonstrated that test-time sensor control improves model robustness more effectively than scaling up the model size or augmenting training data [1, 2].
-However, these efforts have been limited to unimodal RGB tasks.
-Importantly, no prior work has examined multimodal (RGB and depth) sensor control or provided a benchmark that physical sweeps multimodal sensor parameters with ground-truth 6D pose.
-Consequently, the effectiveness of adaptive multimodal sensing and its interplay with traditional data-centric training remain an open question.
 To bridge this gap, we introduce SenseShift6D, a novel RGB-D benchmark for 6D pose estimation that physically sweeps key sensor and lighting parameters.
-Using an Intel RealSense RGB-D camera in a controlled darkroom setup, we systematically vary 13 exposure levels, 9 gain settings, 4 depth-capture modes, and 5 illumination conditions across three representative household objects: a textured spray can, a cylindrical Pringles tube, and a glossy tin case.
-This protocol produces a dataset of 101.9k RGB and 10k depth frames, which can provide 1,380 unique sensor-lighting permutations per object pose (as illustrated in Figures 1 and 2), each annotated via 6D ground-truth via ArUco calibration and manual refinement (with $\sim$3 mm errors).
-For controlled study, experiments in this paper are conducted on the four subsets of the entire dataset: Train-Def, recorded under default camera and illumination settings, and Train-Var, covering a comprehensive multimodal parameter grid (8 exposure, 5 gain, auto-exposure, 4 depth-capture settings with 5 lighting conditions).
+Using an Intel RealSense RGB-D camera in a controlled darkroom setup, we systematically vary 13 exposure levels, 9 gain settings, 4 depth-capture modes, and 5 illumination conditions across five representative household objects: a spray bottle, a Pringles tube, a tincase, a sandwich replica, and a vertical mouse.
+This protocol produces a dataset of 166.4k RGB and 16.7k depth frames, which can provide 1,380 unique sensor-lighting permutations per object pose (Figures 1 and 2), and 665.9k paired RGB-D scenes—to our knowledge, the largest of its kind.
+Each scene is annotated via 6D ground-truth via ArUco calibration and manual refinement (with $\sim$2 mm errors).
+For controlled study, experiments in this paper are conducted on the four subsets of the entire dataset: Train-Def, captured under default camera and illumination settings, and Train-Var, covering a comprehensive multimodal parameter grid (8 exposure, 5 gain, auto-exposure, 1 depth-capture settings with 5 lighting conditions).
 Correspondingly, we construct two test splits (Test-Def and Test-Var) without overlap with the training configurations.
@@
   • Sensor-aware multimodal benchmarking: We present the first 6D object-pose benchmark that systematically and physically varies multimodal sensor parameters across frames, enabling fine-grained analysis of sensor-dependent performance degradation.
-  • Robustness through test-time sensor control: We demonstrate that adjusting sensor parameters during inference under environmental variations improves 6D pose estimation accuracy by an average of 11.4% compared to standard auto-exposure with conventional digital augmentation.
-This indicates that sensor-level tuning can better account for the real-world variability faced during deployment.
-Moreover, sensor control proves to be a data-efficient approach to improving robustness, achieving comparable generalization with significantly less training data.
-  • Potential of multimodal sensor adaptation: While adapting either RGB or depth sensors individually already yields noticeable improvements, jointly optimizing both RGB and depth parameters at test-time provides even greater gains.
-This emphasizes the value of studying multimodal sensor interactions and highlights new directions for advancing the field of 6D pose estimation.
+  • New Axes of Generalization: We reveal that environmental illumination and sensor settings constitute an orthogonal axis of generalization that current approaches do not address: (i) pretrained, generalizable models targeting unseen objects, occlusion, and clutter still degrade markedly under these shifts; and (2) instance-level models trained with standard augmentations likewise degrade even for known objects.
+  • Robustness via multimodal test-time sensor control: We demonstrate that test-time sensor control substantially improves performance without model retraining.
+While single-modal control is already effective (e.g., +12.3 percentage points (pp) on pretrained models), multimodal control delivers further gains (e.g., +19.5 pp on the same models), highlighting the synergy from jointly adapting RGB and depth sensors.
 By providing a structured testbed for sensor-aware object 6D pose estimation, SenseShift6D lays the groundwork for more robust, adaptive perception systems in dynamic real-world environments.
@@
 Table 1: Comparison of 6D pose estimation datasets in terms of sensor and lighting variation.
-Dataset Environmental and sensor variations Objects RGB frames Depth frames
+Dataset Environmental and sensor variations Objects RGB Depth RGB-D
 Illumination Exposure Gain Depth
-LM (ACCV ’12) [7] $\times$ $\times$ $\times$ $\times$ 15 18.2k 18.2k
-LM-O (ECCV ’14) [3] $\times$ $\times$ $\times$ $\times$ 8 1.2k 1.2k
-T-LESS (WACV ’17) [10] $\times$ $\times$ $\times$ $\times$ 30 49k 49k
-YCB-V (RSS ’18) [27] $\times$ $\times$ $\times$ $\times$ 21 133k 133k
-TUD-L (ECCV ’18) [11] 8 locations $\times$ $\times$ $\times$ 3 62k 62k
-HB (ICCVW ’19) [15] 2 locations $\times$ $\times$ $\times$ 33 34.8k 34.8k
-HOPE (IROS ’22) [23] 5 locations $\times$ $\times$ $\times$ 28 238 238
-IPD (CVPR ’24) [14] 3 locations 4 levels $\times$ $\times$ 20 30k 30k
-SenseShift6D (Ours) 5 levels 13 levels 9 levels 4 modes 3 101.9k 10k
+LM (Hinterstoisser et al., 2012a) $\times$ $\times$ $\times$ $\times$ 15 18.2k 18.2k 18.2k
+LM-O (Brachmann et al., 2014) $\times$ $\times$ $\times$ $\times$ 8 1.2k 1.2k 1.2k
+T-LESS (Hodan et al., 2017) $\times$ $\times$ $\times$ $\times$ 30 49k 49k 49k
+YCB-V (Xiang et al., 2017) $\times$ $\times$ $\times$ $\times$ 21 133k 133k 133k
+TUD-L (Hodan et al., 2018b) 8 options $\times$ $\times$ $\times$ 3 62k 62k 62k
+HB (Kaskman et al., 2019) 2 options $\times$ $\times$ $\times$ 33 34.8k 34.8k 34.8k
+HOPE (Sundermeyer et al., 2022) 5 options $\times$ $\times$ $\times$ 28 238 238 238
+IPD (Kalra et al., 2024) 3 options 4 levels $\times$ $\times$ 20 30k 30k 30k
+SenseShift6D (Ours) 5 levels 13 levels 9 levels 4 modes 5 166.4k 16.7k 665.9k
 2.1 6D Pose Estimation Benchmark
 Classic datasets.
-LineMOD (LM) [7] and its occlusion subset (LM-O) [3] pioneered RGB-D pose estimation benchmarks in controlled tabletop setups.
-Subsequent benchmarks such as T-LESS [10], which provided a large collection of texture-less industrial objects, and YCB-V [27], which offered cluttered scenes with temporal sequences, expanded the scope of evaluation.
+LineMOD (LM) (Hinterstoisser et al., 2012a) and its occlusion subset (LM-O) (Brachmann et al., 2014) pioneered RGB-D pose estimation benchmarks in controlled tabletop setups.
+Subsequent benchmarks such as T-LESS (Hodan et al., 2017), which provided a large collection of texture-less industrial objects, and YCB-V (Xiang et al., 2017), which offered cluttered scenes with temporal sequences, expanded the scope of evaluation.
 While these benchmarks provide robust evaluation regarding variations in objects and their poses, all were captured under fixed illumination and camera settings.
 Initial exploration of environmental and sensor variation.
-TUD-L [11], HomebrewedDB (HB) [15], and HOPE [23] introduced variations in illumination conditions but retained fixed camera settings with auto-exposure enabled.
-The recent Industrial Parts Dataset (IPD) [14] was the first benchmark to vary camera parameters, capturing four exposure levels under multiple lighting conditions.
+TUD-L (Hodan et al., 2018b), HomebrewedDB (HB) (Kaskman et al., 2019), and HOPE (Sundermeyer et al., 2022) introduced variations in illumination conditions but retained fixed camera settings with auto-exposure enabled.
+The recent Industrial Parts Dataset (IPD) (Kalra et al., 2024) was the first benchmark to vary camera parameters, capturing four exposure levels under multiple lighting conditions.
 However, IPD focuses on texture-less industrial parts and does not explore variations in RGB gain or depth-sensor modes, thus limiting its suitability for adaptive-sensing research.
@@
 2.2 Adaptive Sensor Control Benchmark
-Recent studies have explored test-time adaptation of camera sensor parameters to improve downstream vision performance.
-ImageNet-ES [1] and ImageNet-ES-Diverse [2] captured real-world scenes using a physical camera, systematically varying aperture, shutter speed, and ISO across multiple lighting conditions for image classification tasks.
+Prior studies (Goudreault et al., 2023; Paul et al., 2022) have explored test-time adaptation of camera parameters to improve downstream vision performance, but these studies lacked explicit benchmarks in real-world.
+Recently, ImageNet-ES (Baek et al., 2024) and ImageNet-ES-Diverse (Baek et al., 2025) captured real-world scenes using a physical camera, systematically varying aperture, shutter speed, and ISO across multiple lighting conditions for image classification tasks.
 These benchmarks demonstrated that physical variation cannot be replicated by synthetic augmentations and dynamically selecting optimal camera parameters at test time can significantly increase the performance of lightweight models, enabling accuracy competitive with larger foundation models without additional training.
@@
 Our SenseShift6D addresses this gap by providing the first realistic testbed for evaluating multimodal, per-frame sensor-control policies alongside conventional data-centric training baselines.
+Table 2: SenseShift6D Configurations for Data split.
+[…]
+Train-Def
+50
+Auto (1 RGB options,
+1 depth option)
+Auto
+Auto
+Default
+RGB: 629, Depth: 629
+1 RGB-D scene/pose
+Train-Var
+5, 25,
+50, 75, 100
+Auto + Manual
+(41 RGB options,
+1 depth options)
+Auto, 2, 4, 19,
+78, 312, 1250,
+5000, 10000
+Auto, 0, 32,
+64, 96, 128
+Default
+RGB: 128.9k, Depth: 3.1k
+225 RGB-D scenes/pose
+Test-Def
+50
+Auto (1 RGB option,
+1 depth option)
+Auto
+Auto
+Default
+RGB: 208, Depth: 208
+1 RGB-D scenes/pose
+Test-Var
+5, 25,
+50, 75, 100
+Auto + Manual
+(21 RGB options,
+4 Depth options)
+Auto, 9, 39,
+156, 625, 2500
+Auto, 16, 48,
+80, 112
+Default,
+High Accuracy,
+High Density,
+Medium Density
+RGB: 21.8k, Depth: 4.1k
+480 RGB-D scenes/pose
 3 SenseShift6D
-Table 2: SenseShift6D Configurations for Data split.
-Data split Brightness (%) Sensor configuration Exposure ($\mu$s) Gain Depth-capture mode Captured images
-Train-Def 50 Auto (1 RGB option, 1 depth option) Auto Auto Default RGB: 385, Depth: 385 1 RGB-D scene/pose
-Train-Var 5, 25, 50, 75, 100 Auto + Manual (41 RGB options, 1 depth options) Auto, 2, 4, 19, 78, 312, 1250, 5000, 10000 Auto, 0, 32, 64, 96, 128 Default RGB: 80.0k, Depth: 1.9k 205 RGB-D scenes/pose
-Test-Def 50 Auto (1 RGB option, 1 depth option) Auto Auto Default RGB: 128, Depth: 128 1 RGB-D scenes/pose
-Test-Var 5, 25, 50, 75, 100 Auto + Manual (21 RGB options, 4 Depth options) Auto, 9, 39, 156, 625, 2500 Auto, 16, 48, 80, 112 Default, High Accuracy, High Density, Medium Density RGB: 13k, Depth: 2.5k 420 RGB-D scenes/pose
-Figure 3: Three household objects collected in SenseShift6D.
+Figure 3: Five household objects collected in SenseShift6D.
 We introduce SenseShift6D, the first RGB-D benchmark that physically sweeps camera and lighting parameters, enabling rigorous evaluation of 6D object pose estimators under realistic environmental and sensor dynamics.
-Using an Intel RealSense D455 mounted in a darkroom, we systematically capture three representative household objects – a textured spray can, a cylindrical Pringles tube, and a glossy tin case (Figure 3) – under auto-exposure, 13 RGB-exposure times, 9 RGB-gain settings, 4 depth-capture presets, and 5 illumination levels (details in Table 2).
-Our dataset consists of 101.9k RGB and 10k depth frames, which can yield 1,380 unique sensor-lighting permutations per object pose.
+Using an Intel RealSense D455 mounted in a darkroom, we systematically capture five representative household objects—spray, Pringles, tincase, sandwich, and mouse—selected to reflect diverse combinations of diverse RGB and geometric features (Figure 3), under auto-exposure, 13 RGB-exposure times, 9 RGB-gain settings, 4 depth-capture presets, and 5 illumination levels (details in Table 2).
+Our dataset consists of 166.4k RGB and 16.7k depth frames, which can yield 1,380 unique sensor-lighting permutations per object pose and 665.9k RGB-D combined scenes, the largest among the current benchamrks to our knowledge.
 Accurate ground truth is obtained via a rigidly mounted ChArUco board and careful hand-eye calibration.
@@
   • Train-Def: Contains only the default sensor and lighting configuration (1 RGB-D scene/pose)
-  • Train-Var: Expands Train-Def, covering (8 exposures $\times$ 5 gains + 1 auto-exposure) $\times$ 1 depth modes $\times$ 5 brightness levels (205 RGB-D scenes/pose).
+  • Train-Var: Expands Train-Def, covering (8 exposures $\times$ 5 gains + 5 auto-exposure) $\times$ 1 depth modes $\times$ 5 brightness levels (225 RGB-D scenes/pose).
   • Test-Def: Includes the default sensor and lighting configuration (1 RGB-D scenes/pose)
-  • Test-Var: Expands Test-Def, featuring (5 exposures $\times$ 4 gains + 1 auto-exposure) configurations, spanning the same depth-mode and illumination range (420 RGB-D scenes/pose).
-As all variations are physically captured rather than digitally simulated, the resulting images inherently retain authentic artifacts, such as non-linear color shifts, photon-shot noise, and depth-mode-specific errors, that synthetic augmentation methods cannot reproduce [1].
+  • Test-Var: Expands Test-Def, featuring (5 exposures $\times$ 4 gains + 4 auto-exposure) configurations, spanning the same depth-mode and illumination range (480 RGB-D scenes/pose).
+As all variations are physically captured rather than digitally simulated, the resulting images inherently retain authentic artifacts, such as non-linear color shifts, photon-shot noise, and depth-mode-specific errors, that synthetic augmentation methods cannot reproduce (Baek et al., 2024).
 SenseShift6D therefore offers the first realistic testbed for analyzing environment- and sensor-aware robustness, enabling side-by-side evaluation of per-frame adaptive-sensing policies against traditional data-centric training strategies.
@@
 The Intel RealSense D455 camera permits manual control of two key RGB sensor parameters: exposure, determining the duration of sensor exposure to incoming light, and gain, which amplifies the captured signal strength.
-  • Exposure grid.
-The D455 supports exposure settings ranging from 1 $\mu$s to 10,000 $\mu$s.
-Empirical tests indicated that exposure times below 1,000 $\mu$s yield image quality comparable to the camera’s auto-exposure setting in well-lit environments.
-Given that this threshold is relatively low within the configurable range, we chose 13 logarithmically spaced exposure values to maximize coverage across practical operating conditions.
+[…]
   • Gain grid.
@@
 This process relies on multiple internal parameters whose individual effects are complex and challenging to fine-tune directly.
-To manage this complexity, we leverage four predefined depth presets provided by Intel, each optimized for different usage scenarios [12]:
+To manage this complexity, we leverage four predefined depth presets provided by Intel, each optimized for different usage scenarios (Hsu et al., 2023):
   • Default: Offers visually clean depth maps with reduced noise and well-defined edges, suitable for general-purpose applications.
@@
 Figure 4: Testbed for data collection.
-Table 3: GT annotation accuracy (mm).
+Table 3: GT annotation accuracy ($m$).
 $\mu_{\delta}$ $\sigma_{\delta}$ $\mu_{|\delta|}$ $med_{|\delta|}$
-RealSense -2.15 4.55 3.75 2.89
+RealSense -0.49 3.30 2.34 1.76
 To capture each scene systematically, we acquired images under all sensor configurations associated with a given data split, while maintaining fixed positions for the object and camera.
@@
 We employed one smart IR remote-control hub to send commands to the turntable and another hub to control three smart bulbs.
-Both hubs received commands from a cloud-based server, which coordinated the entire data-capture capture sequence through an integrated control script (see Figure 4).
+Both hubs received commands from a cloud-based server, which coordinated the entire data-capture sequence through an integrated control script (see Figure 4).
 3.3 Data Processing
@@
 To suppress this effect, we captured successive auto-exposure shots of each scene (5 shots for Train-Def/Var and 4 shots for Test-Def/Var), each preceded by a different gain initialization, and average them to obtain a stable reference image.
-Ground-Truth annotation. For each pose, we used the 50%-brightness, averaged auto-exposure RGB frame and went through the annotation process below, resulting in 2$\sim$3 mm errors.
-  • Object masks: Initially, rough bounding boxes were generated using Grounding DINO [18], guided by shape-aware textual prompts tailored to each object.
+Ground-Truth annotation. For each pose, we used the 50%-brightness, averaged auto-exposure RGB frame and went through the annotation process below, resulting in about 2 mm errors.
+  • Object masks: Initially, rough bounding boxes were generated using Grounding DINO (Liu et al., 2024), guided by shape-aware textual prompts tailored to each object.
 Boxes were manually reviewed and corrected if needed.
-The refined boxes were passed to FastSAM [29] to generate precise instance-level masks.
+The refined boxes were passed to FastSAM (Zhao et al., 2023) to generate precise instance-level masks.
   • Object poses: An initial camera-to-object transformation was provided by detecting the ChArUco markerboard rigidly attached to the turntable.
-The object poses were then refined precisely in 3D using the annotation utility [6] in the BOP toolkit [9] with depth maps captured in High Density mode.
+The object poses were then refined precisely in 3D using the annotation utility (Gouda et al., 2022) in the BOP toolkit (Hodan & Sundermeyer, 2020) with depth maps captured in High Density mode.
 Finally, all refined poses were re-projected onto the RGB images and visually inspected to confirm alignment accuracy.
-  • Accuracy of GT annotation: Following the protocol in [10, 15], we calculate the difference $\delta=d_{c}-d_{r}$, where $d_{c}$ denotes the captured depth map and $d_{r}$ denotes the rendered depth map using the GT pose.
+  • Accuracy of GT annotation: Following the protocol in (Hodan et al., 2017; Kaskman et al., 2019), we calculate the difference $\delta=d_{c}-d_{r}$, where $d_{c}$ denotes the captured depth map and $d_{r}$ denotes the rendered depth map using the GT pose.
 Differences over 5 cm are treated as outliers and omitted.
@@
 4 Experiments
-To assess how sensor control improves efficiency in 6D pose estimation relative to conventional training and inference approaches, we conducted experiments on several state-of-the-art models.
-4.1 Experiment Setup
 Models.
-We chose three 6D object pose estimation models: ZebraPose [22], GDRNPP [19], and HiPose [17].
-ZebraPose is an RGB-based method that leverages features extracted from a ResNet-34 backbone.
-GDRNPP improves upon its predecessor, GDR-Net  [25], by adopting a ConvNeXt  [20] backbone and modifying training details such as learning rate scheduling and weight decay.
-While GDRNPP includes a depth-based pose refinement module, its inference primarily relies on RGB inputs.
-HiPose is an RGB-D model that uses ConvNeXt as the RGB backbone and RandLA-Net  [13] to extract point cloud features from depth.
-These RGB and depth features are fused through a bidirectional interaction module for pose estimation.
-In our experiments, we excluded the refinement stage for ZebraPose and GDRNPP, while evaluating the full pipeline of HiPose.
-Train setup.
-For training, we first generated 10k photorealistic (PBR) synthetic images per object using BlenderProc [4].
-Each model was trained independently on a per-object basis, following one of the three schemes: (1) PBR + Train-Def (default physical images only), (2) PBR + Train-Def with data augmentation (including Gaussian blur, linear contrast adjustment, brightness enhancement and others – see Appendix A for details), and (3) PBR + Train-Var (full sensor and lighting variation).
-ZebraPose and GDRNPP were trained using an NVIDIA RTX 3090 GPU, while HiPose was trained on an NVIDIA A100 GPU.
-During training, we utilized depth images captured under the default-capture mode.
-Metric.
-We used the standard Average Distance (ADD) metric [8] to evaluate model performance.
-ADD measures the average point-wise distance between model points transformed by the predicted pose and the ground-truth pose:
-$\text{ADD}=\frac{1}{m}\sum_{\mathbf{x}\in\mathcal{M}}\left\|(\mathbf{R}\mathbf{x}+\mathbf{T})-(\tilde{\mathbf{R}}\mathbf{x}+\tilde{\mathbf{T}})\right\|,$ (1)
-where $\mathcal{M}$ denotes the set of 3D model points, $\mathbf{x}\in\mathcal{M}$ represents an individual point on the model, $m$ is the total number of points in the model.
-$\mathbf{R}$ and $\mathbf{T}$ are the ground-truth rotation matrix and translation vector, while $\tilde{\mathbf{R}}$ and $\tilde{\mathbf{T}}$ are their predicted counterparts.
-4.2 Analysis on RGB Sensor Control
-Table 4 evaluates the impact of illumination and RGB sensor variations on 6D pose estimation.
-All reported results utilize the ADD metric, using AR@5, which corresponds to recall at a threshold of 5% of the object diameter.
-We compare three sensor control settings: (1) AE (auto-exposure) represents the conventional baseline with automatic sensor settings; (2) Rand (random exposure and gain) indicates average performance across randomly selected sensor configurations; (3) Oracle represents the best achievable performance through optimal sensor control – measuring the probability that at least one sensor-controlled configuration enables correct pose prediction for a given scene.
-Table 4: Performance of various sensor control methods under different training and test settings of SenseShift6D.
-ZebraPose [22] GDRNPP [19] HiPose [17]
+To quantify the effect of illumination and sensor variations—and the gains from test‑time sensor control—relative to conventional training–inference paradigms (large‑scale pretraining and standard supervised pipelines), we evaluate several state‑of‑the‑art methods:
+  • Pretrained, generalizable models (unseen objects; RGB-D): GigaPose (Nguyen et al., 2024), SAM-6D (Lin et al., 2024a), and FoundationPose (Wen et al., 2024).
+GigaPose and SAM-6D were pretrained on the large-scale MegaPose (Labbé et al., 2022), which contains objects from ShapeNet (Chang et al., 2015) and GSO (Downs et al., 2022).
+FoundationPose uses a custom data-generation pipeline with 3D assets from GSO and Objaverse (Deitke et al., 2023).
+  • Instance-level models (known objects): ZebraPose (Su et al., 2022), an RGB-based method with a ResNet-34 backbone; GDRNPP (Liu et al., 2025), which builds on GDR-Net (Wang et al., 2021) using a ConvNeXt (Liu et al., 2022) backbone and depth-based refinement module; and HiPose (Lin et al., 2024b), an RGB-D approach that fuses ConvNeXt-based RGB features with RandLA-Net (Hu et al., 2020) depth features.
+These models are trained on SenseShift6D.
+Sensor control.
+We consider five regimes to evaluate the impact of sensor control: (1) Baseline: default settings for both RGB (e,g, auto-exposure) and depth sensors; (2) Depth-Only: Oracle selection for depth presets and default auto-exposure for RGB; (3) RGB-Only: Oracle selection for RGB exposure and gain, with the Default mode for the depth sensor; (4) Oracle-Fixed: a single best-performing fixed RGB-D configuration, applied to all test scenes; (5) Oracle-Dynamic: per-scene Oracle selection over both RGB and depth.
+Here, Oracle represents an upper bound obtained by evaluating the full sensor-configuration grid and counting a scene as correct if any configuration yields a correct prediction.
+Metrics.
+We use the standard Average Distance (ADD) metric (Hinterstoisser et al., 2012b), which measures the average point-wise distance between model points transformed by the predicted pose and the ground-truth pose.
+Based on this metric, we report two evaluation measures: AUC@[0:0.1], which corresponds to the area under the ADD recall curve up to a threshold of 10% of the object diameter, and AR@5, which denotes the recall at a threshold of 5% of the object diameter.
+Due to space limits, additional metrics—average recall of VSD, MSSD and MSPD (Hodan et al., 2018a)—are reported in Appendix D.
+4.1 Analysis on Pretrained, Generalizable Pose Estimation Models
+Generalizable 6D pose estimators are designed to cope with clutter, occlusion, and novel objects by leveraging large‑scale pretraining.
+SenseShift6D removes these scene‑level confounders—single object, clean background, no occlusion—which isolates the effect of environmental and sensor shifts.
+Evaluating GigaPose, SAM‑6D, and FoundationPose on SenseShift6D thus tests whether robustness acquired through pretraining transfers to an orthogonal and underexplored axis: variations in illumination and RGB/depth sensor parameters that are largely absent from existing benchmarks.
+Table 4 presents the impact of illumination and sensor variations on 6D pose estimation, evaluated with ground-truth object segmentations.
+Given that GigaPose and FoundationPose assume ground-truth masks, while SAM-6D includes its own image segmentation model (ISM), we also provide the results of SAM-6D’s full pipeline in Appendix C for completeness.
+Table 4: AUC@[0:0.1] performance of various sensor control methods for unseen object pose estimation models: GigaPose, SAM-6D, FoundationPose.
+Model Object
+Baseline
+RGB: Auto
+Depth: Default
+Depth-Only
+RGB: Auto
+Depth: Oracle
+RGB-Only
+RGB: Oracle
+Depth: Default
+Oracle-Fixed
+Best Fixed Param.
+Oracle-Dynamic
+RGB: Oracle
+Depth: Oracle
+GigaPose
+ (Nguyen et al., 2024)
+Spray 63.94 72.41 (+8.47) 76.42 (+12.48) 67.98 (+4.04) 81.24 (+17.30)
+Pringles 20.91 46.74 (+25.83) 46.13 (+25.22) 30.82 (+9.91) 70.54 (+49.63)
+Tincase 64.49 72.67 (+8.18) 69.32 (+4.83) 65.51 (+1.02) 76.41 (+11.92)
+Sandwich 64.94 71.92 (+6.98) 76.09 (+11.15) 67.90 (+2.96) 83.88 (+18.94)
+Mouse 56.62 67.40 (+10.78) 67.02 (+10.40) 53.46 (-3.16) 76.71 (+20.09)
+Overall 54.18 66.23 (+12.05) 67.00 (+12.82) 57.13 (+2.95) 77.76 (+23.58)
+SAM-6D
+ (Lin et al., 2024a)
+Spray 79.49 89.41 (+9.92) 87.73 (+8.24) 81.22 (+1.73) 90.80 (+11.31)
+Pringles 63.48 79.17 (+15.69) 80.86 (+17.38) 64.54 (+1.06) 83.98 (+20.50)
+Tincase 72.26 83.02 (+10.76) 85.49 (+13.23) 68.85 (-3.41) 86.53 (+14.27)
+Sandwich 78.39 80.85 (+2.46) 81.59 (+3.20) 78.38 (-0.01) 83.03 (+4.64)
+Mouse 53.82 68.24 (+14.42) 74.63 (+20.81) 57.56 (+3.74) 77.61 (+23.79)
+Overall 69.49 80.14 (+10.65) 82.06 (+12.57) 70.11 (+0.62) 84.39 (+14.90)
+FoundationPose
+ (Wen et al., 2024)
+Spray 84.63 86.52 (+1.89) 87.60 (+2.97) 83.96 (-0.67) 89.01 (+4.38)
+Pringles 34.26 50.72 (+16.46) 50.78 (+16.52) 37.73 (+3.47) 74.69 (+40.43)
+Tincase 39.55 58.41 (+18.86) 70.03 (+30.48) 38.25 (-1.30) 80.95 (+41.40)
+Sandwich 83.17 86.40 (+3.23) 85.12 (+1.95) 83.14 (-0.03) 88.11 (+4.94)
+Mouse 70.22 74.34 (+4.12) 75.86 (+5.64) 70.69 (+0.47) 79.98 (+9.76)
+Overall 62.37 71.28 (+8.91) 73.88 (+11.51) 62.76 (+0.39) 82.55 (+20.18)
+Figure 5: Overall standard deviation of AUC computed across all brightness and scenes evaluated under Baseline, Oracle-Fixed, Oracle-Dynamic.
+Default, Fixed, and Adaptive Multimodal Sensor Configurations.
+Baseline (auto‑exposure RGB + Default depth) yields the lowest accuracy among the five regimes.
+Controlling a single modality already helps: RGB‑Only improves by +12.3 pp or more across all models, indicating that factory auto‑exposure—tuned for human perception—is not model‑optimal.
+Depth‑Only also improves performance by +10.5 pp or more, showing that the default depth preset is likewise suboptimal.
+Furthermore, Oracle-Dynamic (per-scene RGB-and-depth selection) achieves the best results, with +19.5 pp over Baseline and +7.2 pp over the second-best RGB-Only, demonstrating complementary cues and synergy when both sensors are adapted jointly.
+In contrast, Oracle-Fixed (a single global configuration) provides only +1.3 pp on average—and can even underperform Baseline (e.g., SAM-6D on Tincase)—underscoring that adaptive, per-scene control is required.
+Crucially, all gains are obtained without any model retraining, highlighting the practical potential of model-specific sensor adaptation at test time.
+These results motivate deeper study of multimodal test-time sensor control and of which RGB-D combinations most benefit each model.
+Impact on Generalization and Stability.
+Under Baseline, we observe substantial object‑wise disparity: the average gap between the best‑ and worst‑performing objects is 40.0 pp.
+Such gaps arise when objects lack distinctive cues for pose disambiguation or deviate geometrically from pretraining distributions, indicating that current generalizable methods still struggle on some novel objects.
+In contrast, Oracle-Dynamic markedly reduces this disparity to 13.6 pp.
+Moreover, Figure 5 shows that per‑scene sensor control reduces scene‑level variability under the same object.
+The overall AUC standard deviation decreases by 8.8 pp relative to Baseline, with GigaPose on the Sandwich showing the largest reduction of 24.2 pp.
+This indicates higher accuracy and greater stability across diverse scenes.
+Together, these findings demonstrate that adapting inputs via sensor control enhances robustness precisely where large‑scale pretrained models tend to fail.
+While prior advances have emphasized clutter, occlusion, and category novelty, our results reveal that environmental and sensor-setting variation constitute a distinct—and underexplored—axis of difficulty warranting dedicated investigation.
+Qualitative examples of pretrained, generalizable models are presented in Appendix E.1.
+4.2 Analysis on Instance-Level Pose Estimation Models
+Instance‑level methods follow the standard train–test setup in which the training and test sets share the same object domain.
+Building on Section 4.1, our goal here is twofold: (i) to test whether environmental and sensor variations remain challenging once the unseen‑object factor is removed, and (ii) to assess whether test‑time sensor control remains effective even when the training data already includes various augmentations or explicit physical sensor and illumination variation.
+For training, we generated 10k photorealistic (PBR) synthetic images per object using BlenderProc (Denninger et al., 2023).
+Each model was trained independently on a per-object basis under three schemes: (1) PBR + Train-Def (default physical images only), (2) PBR + Train-Def with augmentation (including Gaussian blur, linear contrast adjustment, brightness enhancement and others; see Appendix A), and (3) PBR + Train-Var (full sensor and lighting variation).
+In all experiments, we excluded refinement modules for ZebraPose and GDRNPP, and used the full pipeline for HiPose.
+Since HiPose is the only RGB-D method (ZebraPose and GDRNPP are RGB-only), we report results in two parts.
+Table 5 evaluates the impact of illumination and RGB sensor variations (with the default depth preset) under three sensor-control settings: Auto (auto-exposure), Rand (average performance across randomly selected exposure-gain pairs), and Oracle (per-scene upper bound over the full RGB grid).
+Table 6 evaluates HiPose (training scheme 1) on Test-Var under the multimodal regimes introduced in Section 4.1.
+Table 5: AR@5 performance of various sensor control methods under different training and test settings of SenseShift6D for instance-level pose estimation models: ZebraPose, GDRNPP, HiPose.
+ZebraPose (Su et al., 2022) GDRNPP (Liu et al., 2025) HiPose (Lin et al., 2024b)
 Test-Def Test-Var Test-Def Test-Var Test-Def Test-Var
-Object Train AE AE Rand Oracle AE AE Rand Oracle AE AE Rand Oracle
-Spray PBR + Train-Def 100.0 96.09 52.26 99.51 100.0 96.59 52.02 99.02 100.0 98.73 60.00 100.0
+Object Train Auto Auto Rand Oracle Auto Auto Rand Oracle Auto Auto Rand Oracle
+Spray PBR + Train-Def 100.0 96.09 52.26 99.51 100.0 96.59 52.02 99.02 100.0 98.73 59.98 100.0
 PBR + Train-Def w/ Aug 100.0 98.04 68.46 100.0 100.0 98.54 71.88 100.0 100.0 100.0 89.61 100.0
@@
 PBR + Train-Def w/ Aug 73.80 66.18 40.92 89.99 80.95 79.05 49.17 89.52 97.62 97.81 72.81 100.0
-PBR + Train-Var 90.47 89.04 68.35 95.23 90.48 90.004 72.22 97.14 97.62 97.62 86.86 98.10
-Tincase PBR + Train-Def 95.55 84.43 44.15 96.00 82.22 75.11 43.29 90.67 73.33 60.36 40.27 77.78
-PBR + Train-Def w/ Aug 95.55 88.88 53.37 99.10 73.33 69.78 48.85 94.67 75.56 72.44 51.09 76.00
-PBR + Train-Var 97.77 96.88 73.55 99.55 95.56 95.11 77.05 97.78 75.56 73.87 73.09 75.11
-Overall PBR + Train-Def 93.75 85.41 45.09 96.75 72.60 77.53 43.65 92.91 91.11 85.03 53.00 92.59
-PBR + Train-Def w/ Aug 89.78 84.37 54.25 96.36 79.20 82.45 56.63 94.73 91.06 90.08 71.17 92.00
-PBR + Train-Var 95.26 94.33 73.56 97.44 94.53 94.23 78.40 97.82 91.06 90.50 86.65 91.07
-In Table 4, AE consistently exhibits lower performance on Test-Var compared to Test-Def (captured at the same 50% brightness level as Train-Def), regardless of the model or training strategy.
-This clearly illustrates the necessity of evaluating model robustness under realistic illumination shifts.
-Applying digital augmentation during training generally improves AE performance on Test-Var, yet adversely affects certain cases (e.g., ZebraPose for Pringles and GDRNPP for Tincase), suggesting that augmentation alone may not reliably enhance robustness.
-In contrast, training models on physically collected data (Train-Var, which is 3.5 times larger than Train-Def) consistently yields greater performance gains compared to digital augmentation, highlighting the inherent limitations of synthetic simulation in accurately replicating real-world environmental variations.
-Given that physically collecting large-scale, diverse datasets is often challenging in practice, this performance gap underscores the need for an efficient test-time adaptation methods to effectively address sensor and environmental shifts.
-Model-friendly sensor control vs.
-Traditional auto-exposure.
-Within Test-Var, AE significantly outperforms Rand, showing the critical role of intentional sensor control in managing environmental variability.
-Furthermore, Oracle consistently surpasses AE across all models and training strategies, verifying that the auto-exposure mechanism – typically optimized for human visual perception – is not necessarily optimal for model performance.
-Thus, substantial potential exists for model-specific sensor adaptation at test time, without modifying model architectures or training schemes.
-Notably, Oracle achieves over 13.4% average performance improvement compared to AE for RGB-based models (ZebraPose and GDRNPP).
-Test-time sensor control vs.
-Traditional generalization approaches.
-Oracle without augmentation (i.e., the first training scheme) consistently surpasses AE with augmentation, and even outperforms AE trained on Train-Var in most scenarios.
-In the worst case (GDRNPP), Oracle trained on Train-Def remains comparable to AE trained on Train-Var, showing less than a 2% accuracy gap.
-These superior results, relative to traditional generalization strategies, such as augmentation or manual scaling training data, demonstrate the substantial potential of test-time sensor control for 6D pose estimation in dynamic real-world environments.
-Consequently, mitigating covariate shifts during inference through sensor optimization may be more effective, or complementary, compared to exclusively focusing on model generalization to all possible covariate variations, as currently pursued by foundational-model approaches.
-4.3 Analysis on Multimodal RGB-D Sensor Control
-To comprehensively analyze the impact of multimodal sensor control under environmental variations, we train HiPose (an RGB-D model) following the first training scheme (PBR + Train-Def, captured at 50% brightness level) and evaluate it on the Test-Var split.
-To specifically assess high-precision pose estimation, we report AUC@[0:0.1], which measures the area under the ADD recall curve within 10% of the object diameter.
-We consider five conditions: (1) Baseline utilizes the default settings for both RGB and depth sensors; (2) Depth-Only applies Oracle selection for the depth sensor, while maintaining auto-exposure settings for RGB; (3) RGB-Only applies Oracle selection for the RGB sensor, while using the Default mode for the depth sensor; (4) Oracle-Fixed selects the single best-performing fixed RGB-D configuration and applies it across all test scenes; (5) Oracle-Dynamic applies Oracle selection to both RGB and depth sensors for each scene.
-Table 5: AUC performance of various sensor control methods for HiPose, trained on PBR + Train-Def and tested on Test-Var.
-Object Baseline RGB: Auto Depth: Default Depth-Only RGB: Auto Depth: Oracle RGB-Only RGB: Oracle Depth: Default Oracle-Fixed Best Fixed Param. Oracle-Dynamic RGB: Oracle Depth: Oracle
-Spray 86.80 88.04(+1.24) 88.87(+2.07) 86.80 89.92(+3.12)
-Pringles 73.65 76.10(+2.45) 78.92(+5.27) 74.68(+1.03) 81.42(+7.77)
-Tincase 56.07 58.63(+2.56) 63.60(+7.53) 56.10(+0.03) 66.92 (+10.85)
-Overall 72.17 74.25(+2.08) 77.13(+4.95) 72.53(+0.53) 79.42(+7.24)
-Table 5 shows that all four sensor-control methods outperform Baseline across every scenario.
-Among these methods, Oracle-Fixed yields the smallest improvement, offering only marginal gains over Baseline despite controlling multimodal sensor parameters.
-This result highlights the impracticality of relying on a single, fixed sensor configuration to handle diverse environmental variations; instead, sensor parameters must be adapted dynamically on a per-scene basis.
-Complementary role of RGB and Depth.
-Interestingly, Depth-Only achieves a notable AUC improvement of 2.08% compared to Baseline, demonstrating consistent benefits across all objects.
-Even though the model was trained exclusively with the Default depth-capture mode – suggesting that employing alternative modes at inference might introduce covariate shifts – allowing flexible depth-mode selection still substantially enhances pose estimation performance.
-RGB-Only further improves performance relative to Depth-Only, particularly for symmetric objects (Pringles and Tincase).
-This suggests that RGB surface information may provide crucial cues for accurately estimating the poses of symmetric objects.
-Multimodal joint control vs.
-Unimodal control.
-Oracle-Dynamic achieves the best overall performance, improving AUC by 7.24% over Baseline and by 2.95% over the second-best RGB-Only method.
-Notably, Oracle-Dynamic exhibits the most significant gain for Tincase (10.85% above Baseline), the most challenging object in our dataset.
-These results indicate that RGB and depth modalities provide complementary information, and that jointly adapting both sensors creates synergy exceeding that achievable by independently adapting individual modalities.
-Thus, our findings highlight the importance of deeper investigations into multimodal sensor adaptation strategies and understanding which combinations of RGB-D inputs are most beneficial for model performance.
-4.4 Qualitative Analysis
-Analysis on RGB sensor control.
-Figure 5 illustrates two representative test scenes for GDRNPP, involving the Tincase and Pringles objects, respectively.
-In both cases, AE results in incorrect pose predictions (via GDRNPP), while Oracle (RGB-only) successfully estimates accurate object poses.
-These scenes were captured at the lowest brightness level, representing the most challenging environmental condition.
-These examples highlight that optimal RGB sensor parameters, especially gain values in these cases, vary depending on both object type and object pose, underscoring the necessity of per-scene test-time sensor control.
-Furthermore, the RGB images selected by Oracle and AE appear visually indistinguishable from a human perspective.
-This indicates that the optimal direction for model-friendly sensor control is not always intuitively aligned with human visual perception, emphasizing the importance of sensor control strategies specifically tailored to model performance rather than subjective human assessment.
-Analysis on depth sensor control.
-Figure 6 illustrates two representative test scenes evaluated with HiPose, involving the Tincase and Pringles objects, respectively.
-Both scenes were captured at the lowest brightness level, representing the most challenging environmental conditions.
-While the RGB sensor parameters were fixed, we compared pose estimation performance across different depth-capture modes.
-In both cases, the Default depth mode resulted in incorrect pose predictions, whereas Oracle (Depth-only) successfully produced accurate object poses.
-These examples clearly demonstrate that different depth-capture modes yield varying depth-map qualities, affecting pose estimation outcomes.
-Furthermore, the optimal depth-capture mode depends on the specific object type and scene configuration, underscoring the importance of adaptive, scene-dependent depth-sensor control strategies.
-Figure 5: RGB visualization of pose estimation results under auto exposure and sensor control settings. The red line represents the ground truth pose, and the green line indicates the predicted pose.
-Figure 6: Visualization of pose estimation results showing that depth modes other than the default yield better performance under a fixed RGB configuration (E2500, G16).
-Red lines denote ground truth poses, while green lines represent predicted poses.
+PBR + Train-Var 90.47 89.04 68.35 95.23 90.48 90.00 72.22 97.14 97.62 97.62 86.86 98.10
+Tincase PBR + Train-Def 95.56 79.56 52.53 96.44 91.11 81.33 46.58 92.89 97.78 81.51 53.62 96.44
+PBR + Train-Def w/ Aug 97.78 89.78 65.46 100.0 80.00 82.22 52.89 95.11 97.78 95.64 86.47 97.78
+PBR + Train-Var 98.22 97.60 79.59 100.0 100.0 99.56 82.02 100.0 97.78 97.69 97.76 97.78
+Sandwich PBR + Train-Def 94.87 89.23 48.26 100.0 97.44 88.21 47.03 96.41 94.87 87.69 57.21 94.87
+PBR + Train-Def w/ Aug 92.31 88.72 57.82 98.97 94.87 92.31 63.72 98.97 94.87 93.74 84.38 94.87
+PBR + Train-Var 89.74 89.23 67.15 96.93 97.44 97.44 82.64 100.0 94.87 94.87 94.87 94.87
+Mouse PBR + Train-Def 70.73 61.46 27.37 86.34 95.12 76.10 31.44 88.78 90.24 62.93 43.73 87.81
+PBR + Train-Def w/ Aug 48.78 49.76 32.56 88.29 78.05 72.68 43.29 86.83 90.24 89.07 82.24 90.24
+PBR + Train-Var 46.34 46.34 34.98 73.17 80.49 80.49 66.39 91.22 90.24 89.95 87.17 90.73
+Overall PBR + Train-Def 89.37 80.41 43.86 95.41 90.54 80.64 42.54 93.23 96.56 85.37 54.66 95.82
+PBR + Train-Def w/ Aug 82.09 76.45 50.46 94.74 86.77 84.96 56.19 94.09 96.10 95.25 83.10 96.58
+PBR + Train-Var 84.47 83.86 65.77 92.58 93.19 93.01 77.84 97.38 96.10 96.03 93.33 96.30
+Sensor and Environmental Shift.
+In Table 5, when trained with Train-Def, Auto performance is consistently lower on Test-Var than Test-Def, reflecting sensitivity to illumination shifts.
+In contrast, Oracle on Test‑Var not only surpasses Auto on Test‑Var but often exceeds Auto on Test‑Def, showing that adapting RGB sensor settings at test time both mitigates domain shift and unlocks higher accuracy.
+Table 6 (RGB-D) reveals that while all sensor‑control methods improve over Baseline, depth‑only control provides marginal gains relative to RGB control—unlike the trend in Table 4.
+This is expected because: (i) instance‑level training for HiPose exposes the model to the exact objects, yielding object‑specific optimization that leaves less head‑room for improved depth measurements; (ii) training already uses depth maps from the same domain as Baseline; and (iii) depth sensing is inherently less sensitive to illumination changes than RGB.
+Even with identical objects and backgrounds, however, performance still varies under environmental and sensor changes.
+Table 6: AUC@[0:0.1] performance of various sensor control methods for HiPose, trained on PBR + Train-Def and tested on Test-Var.
+Object
+Baseline
+RGB: Auto
+Depth: Default
+Depth-Only
+RGB: Auto
+Depth: Oracle
+RGB-Only
+RGB: Oracle
+Depth: Default
+Oracle-Fixed
+Best Fixed Param.
+Oracle-Dynamic
+RGB: Oracle
+Depth: Oracle
+Spray 86.92 88.17 (+1.25) 88.79 (+1.87) 87.54 (+0.62) 89.87 (+2.95)
+Pringles 73.40 76.18 (+2.78) 79.08 (+5.68) 77.02 (+3.62) 81.55 (+8.15)
+Tincase 74.78 77.88 (+3.10) 85.30 (+10.52) 83.57 (+8.79) 86.84 (+12.06)
+Sandwich 78.68 81.72 (+3.04) 84.06 (+5.38) 82.90 (+4.22) 86.43 (+7.75)
+Mouse 63.92 66.41 (+2.49) 72.19 (+8.27) 69.52 (+5.60) 75.21 (+11.29)
+Overall 75.54 78.07 (+2.53) 81.88 (+6.34) 80.11 (+4.57) 83.98 (+8.44)
+[…]
+These results demonstrate the substantial potential of test‑time sensor control relative to conventional generalization strategies (augmentation or scaling training data), even for known objects.
+Mitigating covariate shift at inference via sensor optimization can be more effective—or complementary—than exclusively attempting to generalize the model to all possible shifts.
+Qualitative results are provided in Appendix E.2.
 5 Conclusion
-We introduced SenseShift6D, the first sensor-aware, multimodal benchmark for 6D object-pose estimation that physically sweeps RGB exposure, RGB gain, depth-capture mode, and ambient illumination.
-Extensive experiments show that state-of-the-art methods suffer substantial accuracy drops under these realistic sensor shifts.
-Crucially, we demonstrated that adaptive sensor control at test time can recover a large portion of this lost performance – without any model retraining – and that joint RGB + depth control outperforms unimodal strategies.
-These findings highlight sensor control as a promising and lightweight avenue for deploying 6D-pose systems in dynamic, real-world conditions.
+We introduced SenseShift6D, a physically captured RGB‑D benchmark that orthogonally sweeps exposure, gain, depth‑capture mode, and illumination to isolate the impact of sensor and environmental shifts on 6D object pose estimation.
+Across pretrained, generalizable models (GigaPose, SAM‑6D, FoundationPose) and instance‑level models (ZebraPose, GDRNPP, HiPose), we found that factory defaults are not model‑optimal: RGB‑Only control improves accuracy by +12.3 pp or more and Depth‑Only by +10.5 pp or more, while multimodal (RGB+Depth) control achieves the best results (+19.5 pp over Baseline; +7.2 pp over RGB‑Only).
+A single global configuration (Oracle‑Fixed) yields only +1.3 pp on average (and can underperform Baseline), underscoring the need for adaptive, per‑scene control.
+Beyond mean accuracy, per‑scene control reduces object‑wise disparity from 40.0 pp to 13.6 pp and lowers scene‑level variability by 8.8 pp, indicating improved stability as well as accuracy.
+Crucially, these gains require no retraining, positioning test‑time sensor control as a data‑ and compute‑efficient complement to model‑centric scaling.
 Limitations and Future Work.
@@
 The top-left image is the original, and the remaining images are randomly augmented samples generated using the settings summarized in Table 7.
-Appendix B Licenses
-We use open-source components under various licenses for evaluation and data generation.
-In particular, some scripts for photorealistic rendering are adapted from BlenderProc (GPL-3.0), and evaluation pipelines integrate components from GDRNPP (Apache-2.0), ZebraPose (MIT), and HiPose (MIT).
-All third-party code is used in compliance with their respective licenses, with attribution provided as required.
-Our dataset, SenseShift6D, is distributed under the Creative Commons Attribution 4.0 International (CC BY 4.0) license.
-Appendix C Experimental Details
-C.1 Compute Resources
+Appendix B Experimental Details
+B.1 Computing Resources
 All models were trained on a high-performance computing cluster equipped with NVIDIA GPUs and Intel Xeon CPUs.
@@
 Each training node had access to 512GB of RAM.
-C.2 Model Training
+B.2 Pretrained Weights for Unseen Object Pose Estimation Models
+For GigaPose, we used the pretrained weight files ‘gigaPose_v1’ (coarse) and ‘refiner-rgbd-288182519’ (refiner), following instructions from GigaPose GitHub repository.
+For FoundationPose, we used the pretrained weight files ‘2023-10-28-18-33-37’ (refiner) and ‘2024-01-11-20-02-45’ (scorer), following instructions from FoundationPose GitHub repository.
+For SAM-6D, we used the pretrained weight file ‘vit_h’ for SAM and the ‘sam-6d-pem-base.pth’ file for the Pose Estimation Module (PEM), following instructions from SAM-6D Github repository.
+B.3 Model Training for Instance-Level Pose Estimation Models
 All models were trained with a batch size of 32.
@@
 All learning rate schedules and optimizer settings follow the original implementations of each method.
-Table 7: Configurations for various data split.
-Data split Brightness (%) Sensor configuration Exposure ($\mu$s) Gain Depth-capture mode
-Train-B5Def 5 Auto (1 RGB option, 1 depth option) Auto Auto Default
-Train-BADef 5, 25, 50, 75, 100 Auto (1 RGB option, 1 depth option) Auto Auto Default
-Train-B50Var 50 Auto + Manual (41 RGB options, 1 depth options) Auto, 2, 4, 19, 78, 312, 1250, 5000, 10000 Auto, 0, 32, 64, 96, 128 Default
-Test-B5Def 5 Auto (1 RGB option, 1 depth option) Auto Auto Default, High Accuracy, High Density, Medium Density
-Test-B5Var 5 Auto + Manual (21 RGB options, 4 Depth options) Auto, 9, 39, 156, 625, 2500 Auto, 16, 48, 80, 112 Default, High Accuracy, High Density, Medium Density
-Appendix D Additional Evaluation Results
-Table 8 presents the performance of the Tincase in Table 4, categorized by brightness level.
-This suggests a noticeable difference in performance at brightness 5 compared to other brightness levels.
-Therefore, we evaluated SenseShift6D dataset on various data splits, beyond those described in Table 2.
-Each split is detailed in Table 7.
-Table 8: Performance of sensor control across brightness levels for Tincase.
-ZebraPose GDRNPP HiPose
-Brightness Test-Var PBR+ Train-Def PBR+ Train-Def w/ Aug PBR+ Train-Var PBR+ Train-Def PBR+ Train-Def w/ Aug PBR+ Train-Var PBR+ Train-Def PBR+ Train-Def w/ Aug PBR+ Train-Var
-B5 AE 33.33 64.44 95.55 37.78 40.00 93.33 10.67 64.89 71.56
-Rand 9.67 15.56 36.89 13.56 19.44 42.55 10.78 18.67 69.11
-Oracle 80.00 97.78 97.78 88.89 91.11 100.00 75.56 71.11 71.11
-B25 AE 97.78 93.33 95.56 84.44 73.33 95.56 71.11 73.33 73.33
-Rand 46.11 51.11 70.00 43.89 46.22 75.11 38.67 44.56 73.33
-Oracle 100.00 97.78 100.00 91.11 95.56 97.78 77.78 73.33 73.33
-B50 AE 95.56 95.56 97.78 82.22 75.56 95.56 72.44 75.11 73.78
-Rand 56.11 64.67 87.00 55.33 58.00 88.34 53.00 60.89 73.56
-Oracle 100.00 100.00 100.00 91.11 97.78 97.78 77.78 77.78 75.56
-B75 AE 97.78 93.33 97.78 84.44 80.00 95.56 73.78 74.22 76.44
-Rand 54.22 69.44 89.78 50.22 62.78 92.11 46.78 69.44 75.33
-Oracle 100.00 100.00 100.00 91.11 95.56 97.78 80.00 80.00 77.78
-B100 AE 97.78 97.78 97.78 86.67 80.00 95.56 73.78 74.67 74.22
-Rand 54.67 66.11 84.11 53.45 57.78 87.11 52.11 61.89 74.11
-Oracle 100.00 100.00 100.00 91.11 93.33 95.56 77.78 77.78 77.78
-Overall AE 84.43 88.88 96.88 75.11 69.78 95.11 60.36 72.44 73.87
-Rand 44.15 53.37 73.55 42.29 48.85 77.05 40.27 51.09 73.09
-Oracle 96.00 99.10 99.55 90.67 94.67 97.78 77.78 76.00 75.11
-D.1 Trained without Environmental Shift
-We conducted an experiment under the low-light condition of brightness 5, which yielded the lowest performance in brightness-wise evaluation.
-To eliminate environmental shifts, we set both training and testing to brightness 5, replacing the default training condition (brightness 50).
-As shown in Table 9, the oracle with sensor control achieves performance comparable to or even better than training and testing conducted entirely within the same brightness 5 domain, suggesting that adaptive sensor control can mitigate domain gaps without the need for retraining in the target environment.
-Table 9: Comparison of models trained without environmental shift under brightness 5.
-Model Train-B5Def Train-Def
-Test-B5Def AE Test-B5Def AE Test-B5Var Rand Test-B5Var Oracle
-ZebraPose 26.66 33.33 9.66 80.00
-GDRNPP 86.67 37.78 13.56 88.89
-HiPose 35.44 10.67 10.78 75.56
-D.2 Trained on Diverse Sensor and Environmental Settings
-In contrast to Train-Var, we partially incorporated sensor control and environmental variability into the training process.
-Table 10 presents AUC performance at [0:0.1], indicating that including sensor-controlled images at a fixed brightness level (Train-B50Var) can lead to improved generalization across diverse environmental conditions, when compared to using auto exposure images from multiple brightness levels (Train-BADef).
-This indicates that sensor control introduces rich variations which act as implicit regularization against domain shifts.
-Notably, even without additional training, appropriate sensor configuration at test time contributes significantly to model performance, as shown by the oracle performance under the Train-Def setting.
-Table 10: AUC performance under various brightness and sensor configurations for Tincase.
-ZebraPose GDRNPP HiPose
-Test-Def Test-Var Test-Def Test-Var Test-Def Test-Var
-Train AE AE Rand Oracle AE AE Rand Oracle AE AE Rand Oracle
-Train-Def 76.80 68.14 38.20 84.62 67.34 60.93 31.51 77.50 61.58 55.90 34.78 64.22
-Train-BADef 78.73 70.09 37.31 83.96 81.54 80.38 48.83 89.04 61.82 55.36 34.32 63.73
-Train-B50Var 78.91 77.00 56.16 85.58 81.17 79.69 62.03 87.48 61.82 61.54 60.47 63.00
-Train-Var 80.44 79.54 61.91 86.60 79.32 78.64 64.20 85.46 61.73 61.61 60.86 63.13
-Appendix E Additional Qualitative Results on Depth Variations
-Figure 8 visualizes how the object region appears in the depth map under each depth capture mode for all three objects.
-In addition, Figure 9 shows qualitative examples for Tincase, using HiPose trained on Train-Def.
-Given a fixed RGB image with poor visibility, the figure compares predictions and corresponding masked depth maps between the default mode and oracle-selected modes, where the default mode fails across all examples while the oracle modes often produce better results including some successful predictions.
-Figure 8: Visualization of masked object depth maps across different capture modes.
-Figure 9: Comparison of predictions under default and oracle capture modes with a fixed RGB image.
-(Left) Visualized object pose on RGB images: ground truth pose in red, predicted pose in green.
-(Right) Masked depth maps corresponding to each capture mode.
+Appendix C Evaluation of the full SAM-6D pipeline
+Table 8 presents SAM-6D performance based on the Top-30 segmentation results produced by its ISM.
+Table 8: AUC@[0:0.1] performance of SAM-6D with its Image Segmentation Model (ISM).
+Object
+Baseline
+RGB: Auto
+Depth: Default
+Depth-Only
+RGB: Auto
+Depth: Oracle
+RGB-Only
+RGB: Oracle
+Depth: Default
+Oracle-Fixed
+Best Fixed Param.
+Oracle-Dynamic
+RGB: Oracle
+Depth: Oracle
+Spray 81.17 88.93 (+7.76) 89.88 (+8.71) 78.71 (-2.46) 90.66 (+9.49)
+Pringles 61.84 78.71 (+16.87) 79.76 (+17.92) 55.85 (-5.99) 83.14 (+21.30)
+Tincase 63.85 82.13 (+18.28) 82.16 (+18.31) 60.02 (-3.83) 85.42 (+21.57)
+Sandwich 74.66 81.13 (+6.47) 81.62 (+6.96) 71.47 (-3.19) 82.82 (+8.16)
+Mouse 49.16 66.49 (+17.33) 71.46 (+22.30) 47.96 (-1.20) 74.37 (+25.21)
+Overall 66.13 79.48 (+13.35) 80.97 (+14.84) 62.80 (-3.33) 83.28 (+17.15)
+Appendix D Evaluation on BOP metrics
+Table 9 reports the evaluation of unseen object pose estimation models—GigaPose, SAM-6D, and FoundationPose—using the BOP metrics: VSD, MSSD, and MSPD.
+Table 9: Average Reacll of VSD, MSSD, and MSPD on various sensor control methods for unseen object pose estimation models: GigaPose, SAM-6D, FoundationPose.
+Model Object
+Baseline
+RGB: Auto
+Depth: Default
+Depth-Only
+RGB: Auto
+Depth: Oracle
+RGB-Only
+RGB: Oracle
+Depth: Default
+Oracle-Fixed
+Best Fixed Param.
+Oracle-Dynamic
+RGB: Oracle
+Depth: Oracle
+GigaPose
+ (Nguyen et al., 2024)
+Spray 71.42 / 82.68 / 84.15 84.46 / 88.68 / 90.39 83.19 / 94.05 / 94.54 77.76 / 86.15 / 87.71 91.44 / 95.22 / 96.20
+Pringles 79.30 / 68.24 / 84.29 88.88 / 83.14 / 95.33 93.55 / 87.90 / 95.10 84.03 / 75.76 / 90.00 97.72 / 96.05 / 99.43
+Tincase 85.04 / 95.87 / 99.20 90.87 / 99.16 / 99.96 90.72 / 97.16 / 99.60 86.87 / 94.93 / 97.38 96.48 / 99.60 / 100.0
+Sandwich 80.50 / 83.28 / 86.15 88.76 / 87.54 / 87.69 91.89 / 96.31 / 99.18 85.12 / 85.64 / 86.36 98.21 / 99.95 / 100.0
+Mouse 80.26 / 84.44 / 94.34 88.41 / 91.02 / 96.44 86.94 / 91.90 / 98.34 79.34 / 79.76 / 88.34 93.90 / 96.78 / 99.61
+Overall 79.31 / 82.90 / 89.63 88.28 / 89.91 / 93.96 89.26 / 93.46 / 97.35 82.62 / 84.45 / 89.96 95.55 / 97.52 / 99.05
+SAM-6D
+ (Lin et al., 2024a)
+Spray 92.47 / 96.33 / 96.53 94.96 / 98.00 / 98.15 91.00 / 93.66 / 93.66 92.77 / 96.33 / 96.53 97.43 / 99.22 / 99.17
+Pringles 94.38 / 93.08 / 98.56 97.05 / 97.76 / 100.0 95.96 / 96.57 / 98.57 94.32 / 92.43 / 96.81 99.16 / 99.62 / 100.0
+Tincase 97.53 / 88.95 / 90.32 96.47 / 93.33 / 94.36 96.00 / 94.67 / 95.16 97.31 / 82.15 / 83.52 97.33 / 97.20 / 97.82
+Sandwich 88.74 / 94.87 / 94.87 89.42 / 94.87 / 94.92 90.19 / 94.87 / 95.38 88.65 / 94.87 / 94.87 91.32 / 95.44 / 96.31
+Mouse 78.79 / 79.49 / 85.10 82.08 / 84.41 / 89.12 84.92 / 87.06 / 91.42 79.04 / 81.56 / 86.87 88.56 / 89.85 / 93.92
+Overall 90.38 / 90.54 / 93.08 92.00 / 93.68 / 95.31 91.62 / 93.37 / 94.84 90.42 / 89.47 / 91.72 94.76 / 96.27 / 97.44
+FoundationPose
+ (Wen et al., 2024)
+Spray 89.45 / 99.22 / 99.32 92.85 / 99.76 / 99.85 94.27 / 100.0 / 100.0 89.49 / 99.32 / 99.37 97.15 / 100.0 / 100.0
+Pringles 93.98 / 58.29 / 75.29 96.16 / 77.00 / 86.38 96.41 / 81.10 / 89.52 94.14 / 62.43 / 77.38 99.12 / 96.10 / 98.33
+Tincase 97.36 / 44.89 / 49.11 97.60 / 65.64 / 68.09 97.62 / 79.11 / 81.73 96.96 / 43.56 / 47.33 97.75 / 89.69 / 92.09
+Sandwich 94.49 / 95.08 / 96.31 97.26 / 97.59 / 98.97 95.75 / 96.26 / 98.26 95.87 / 97.38 / 97.85 97.96 / 98.41 / 99.49
+Mouse 88.39 / 90.59 / 94.73 91.24 / 92.59 / 97.66 90.72 / 91.95 / 96.68 89.02 / 91.56 / 96.54 93.66 / 94.15 / 98.44
+Overall 92.73 / 77.61 / 82.95 95.02 / 86.52 / 90.19 94.95 / 89.68 / 93.24 93.10 / 78.85 / 83.69 97.13 / 95.67 / 97.67
+Appendix E Qualitative Analysis
+E.1 Qualitative Result on Pretrained, Generalizable Pose Estimation Models
+Figure 789 visualize GT and predicted poses on Baseline and Oracle sensor configuration for GigaPose, SAM-6D and FoundationPose.
+Figure 7: Comparison of predictions under Baseline and Oracle for GigaPose. Visualized object pose on RGB images: ground truth pose in red, predicted pose in green.
+Figure 8: Comparison of predictions under Baseline and Oracle for SAM-6D. Visualized object pose on RGB images: ground truth pose in red, predicted pose in green.
+Figure 9: Comparison of predictions under Baseline and Oracle for FoundationPose. Visualized object pose on RGB images: ground truth pose in red, predicted pose in green.
+E.2 Qualitative Result on Instance-Level Pose Estimation Model
+Figure 10 visualizes GT and predicted poses on Baseline and Oracle sensor configuration for HiPose.
+Figure 10: Comparison of predictions under Baseline and Oracle for HiPose. Visualized object pose on RGB images: ground truth pose in red, predicted pose in green.
+Appendix F Evaluation on RGB-D Based Models Under Different Environments
+Table 10111213 shows performance of RGB-D models—HiPose, GigaPose, SAM-6D, and FoundationPose—under different illumination conditions.
+Table 10: AUC@[0:0.1] performance under diverse environments of various sensor control methods for HiPose.
+Environment Object
+Baseline
+RGB: Auto
+Depth: Default
+Depth-Only
+RGB: Auto
+Depth: Oracle
+RGB-Only
+RGB: Oracle
+Depth: Default
+Oracle-Fixed
+Best Fixed Param.
+Oracle-Dynamic
+RGB: Oracle
+Depth: Oracle
+B5
+Spray 83.17 85.07 (+1.90) 87.10 (+3.93) 86.07 (+2.90) 88.63 (+5.46)
+Pringles 72.40 75.45 (+3.05) 78.60 (+6.19) 77.90 (+5.50) 81.12 (+8.71)
+Tincase 32.96 43.00 (+10.04) 78.80 (+45.84) 76.69 (+43.73) 81.31 (+48.36)
+Sandwich 66.85 69.85 (+3.00) 80.72 (+13.87) 78.82 (+11.97) 83.18 (+16.33)
+Mouse 35.29 39.88 (+4.59) 64.85 (+29.56) 62.93 (+27.63) 69.88 (+34.59)
+Overall 58.13 62.65 (+4.52) 78.01 (+19.88) 76.48 (+18.35) 80.82 (+22.69)
+B25
+Spray 87.56 88.71 (+1.15) 88.61 (+1.05) 87.59 (+0.02) 89.63 (+2.07)
+Pringles 74.12 77.17 (+3.05) 79.14 (+5.02) 77.31 (+3.19) 81.74 (+7.62)
+Tincase 85.20 86.33 (+1.13) 86.91 (+1.71) 85.13 (-0.07) 87.89 (+2.69)
+Sandwich 80.54 83.51 (+2.97) 83.31 (+2.77) 82.97 (+2.44) 86.51 (+5.97)
+Mouse 70.41 73.17 (+2.76) 74.32 (+3.90) 71.54 (+1.12) 77.29 (+6.88)
+Overall 79.57 81.78 (+2.21) 82.46 (+2.89) 80.91 (+1.34) 84.61 (+5.04)
+B50
+Spray 87.93 88.78 (+0.85) 89.32 (+1.39) 88.07 (+0.15) 90.07 (+2.15)
+Pringles 74.24 76.57 (+2.33) 79.69 (+5.45) 76.74 (+2.50) 81.62 (+7.38)
+Tincase 85.11 86.69 (+1.58) 86.91 (+1.80) 85.29 (+0.18) 88.31 (+3.20)
+Sandwich 81.38 84.38 (+3.00) 85.72 (+4.33) 84.03 (+2.64) 87.23 (+5.85)
+Mouse 71.93 73.17 (+1.24) 74.51 (+2.59) 71.63 (-0.29) 77.44 (+5.51)
+Overall 80.12 81.92 (+1.80) 83.23 (+3.11) 81.15 (+1.03) 84.93 (+4.81)
+B75
+Spray 88.02 89.02 (+1.00) 89.51 (+1.49) 88.02 (+0.00) 90.61 (+2.59)
+Pringles 73.31 76.02 (+2.71) 79.12 (+5.81) 76.83 (+3.52) 81.86 (+8.55)
+Tincase 85.27 86.49 (+1.22) 86.80 (+1.53) 85.33 (+0.07) 88.13 (+2.87)
+Sandwich 82.10 85.13 (+3.03) 84.41 (+2.31) 84.38 (+2.28) 87.10 (+5.00)
+Mouse 71.73 73.27 (+1.54) 74.17 (+2.44) 71.24 (-0.49) 76.07 (+4.34)
+Overall 80.09 81.99 (+1.90) 82.80 (+2.71) 81.16 (+1.07) 84.75 (+4.66)
+B100
+Spray 87.93 89.27 (+1.34) 89.39 (+1.46) 87.95 (+0.02) 90.41 (+2.49)
+Pringles 72.90 75.69 (+2.79) 78.86 (+5.95) 76.31 (+3.40) 81.40 (+8.50)
+Tincase 85.38 86.87 (+1.49) 87.09 (+1.71) 85.40 (+0.02) 88.53 (+3.16)
+Sandwich 82.51 85.72 (+3.21) 86.15 (+3.64) 84.31 (+1.80) 88.10 (+5.59)
+Mouse 70.24 72.59 (+2.34) 73.07 (+2.83) 70.24 (+0.00) 75.39 (+5.15)
+Overall 79.79 82.03 (+2.24) 82.91 (+3.12) 80.84 (+1.05) 84.77 (+4.98)
+Table 11: AUC@[0:0.1] performance under diverse environments of various sensor control methods for GigaPose.
+Environment Object
+Baseline
+RGB: Auto
+Depth: Default
+Depth-Only
+RGB: Auto
+Depth: Oracle
+RGB-Only
+RGB: Oracle
+Depth: Default
+Oracle-Fixed
+Best Fixed Param.
+Oracle-Dynamic
+RGB: Oracle
+Depth: Oracle
+B5
+Spray 65.76 73.34 (+7.58) 73.39 (+7.63) 66.51 (+0.75) 78.41 (+12.65)
+Pringles 19.17 44.19 (+25.02) 44.64 (+25.47) 31.33 (+12.16) 66.69 (+47.52)
+Tincase 60.69 71.51 (+10.82) 65.20 (+4.51) 65.89 (+5.20) 75.07 (+14.38)
+Sandwich 59.90 68.82 (+8.92) 69.46 (+9.56) 61.36 (+1.46) 80.28 (+20.38)
+Mouse 48.27 60.44 (+12.17) 63.80 (+15.53) 55.59 (+7.32) 74.29 (+26.02)
+Overall 50.76 63.66 (+12.90) 63.30 (+12.54) 56.14 (+5.38) 74.95 (+24.19)
+B25
+Spray 67.44 73.20 (+5.76) 77.78 (+10.34) 70.49 (+3.05) 81.15 (+13.71)
+Pringles 15.62 49.24 (+33.62) 49.36 (+33.74) 38.17 (+22.55) 73.88 (+58.26)
+Tincase 63.07 72.13 (+9.06) 69.29 (+6.22) 67.29 (+4.22) 75.69 (+12.62)
+Sandwich 68.08 74.18 (+6.10) 77.72 (+9.64) 75.79 (+7.71) 84.59 (+16.51)
+Mouse 61.93 69.71 (+7.78) 68.15 (+6.22) 61.95 (+0.02) 77.00 (+15.07)
+Overall 55.23 67.69 (+12.46) 68.46 (+13.23) 62.74 (+7.51) 78.46 (+23.23)
+B50
+Spray 65.51 73.71 (+8.20) 77.20 (+11.69) 71.10 (+5.59) 83.78 (+18.27)
+Pringles 24.43 45.40 (+20.97) 43.33 (+18.90) 33.50 (+9.07) 70.14 (+45.71)
+Tincase 64.98 72.42 (+7.44) 70.22 (+5.24) 68.31 (+3.33) 76.82 (+11.84)
+Sandwich 63.92 70.46 (+6.54) 77.51 (+13.59) 78.03 (+14.11) 84.38 (+20.46)
+Mouse 56.93 68.93 (+12.00) 67.98 (+11.05) 63.68 (+6.75) 78.20 (+21.27)
+Overall 55.15 66.18 (+11.03) 67.25 (+12.10) 62.92 (+7.77) 78.67 (+23.52)
+B75
+Spray 61.41 71.10 (+9.69) 77.37 (+15.96) 70.68 (+9.27) 81.41 (+20.00)
+Pringles 22.88 47.57 (+24.69) 45.69 (+22.81) 35.12 (+12.24) 72.00 (+49.12)
+Tincase 66.78 73.36 (+6.58) 71.49 (+4.71) 70.78 (+4.00) 77.36 (+10.58)
+Sandwich 66.72 73.00 (+6.28) 77.79 (+11.07) 78.44 (+11.72) 84.95 (+18.23)
+Mouse 55.59 68.66 (+13.07) 67.63 (+12.04) 63.61 (+8.02) 76.93 (+21.34)
+Overall 54.68 66.74 (+12.06) 67.99 (+13.31) 63.73 (+9.05) 78.53 (+23.85)
+B100
+Spray 59.59 70.71 (+11.12) 76.37 (+16.78) 70.71 (+11.12) 81.44 (+21.85)
+Pringles 22.48 47.31 (+24.83) 47.64 (+25.16) 33.12 (+10.64) 70.00 (+47.52)
+Tincase 66.93 73.91 (+6.98) 70.42 (+3.49) 69.38 (+2.45) 77.13 (+10.20)
+Sandwich 66.10 73.15 (+7.05) 77.95 (+11.85) 76.77 (+10.67) 85.21 (+19.11)
+Mouse 60.39 69.29 (+8.90) 67.54 (+7.15) 63.12 (+2.73) 77.15 (+16.76)
+Overall 55.10 66.87 (+11.77) 67.98 (+12.88) 62.62 (+7.52) 78.18 (+23.08)
+Table 12: AUC@[0:0.1] performance under diverse environments of various sensor control methods for SAM-6D.
+Environment Object
+Baseline
+RGB: Auto
+Depth: Default
+Depth-Only
+RGB: Auto
+Depth: Oracle
+RGB-Only
+RGB: Oracle
+Depth: Default
+Oracle-Fixed
+Best Fixed Param.
+Oracle-Dynamic
+RGB: Oracle
+Depth: Oracle
+B5
+Spray 80.41 86.83 (+6.42) 84.27 (+3.86) 82.66 (+2.25) 89.27 (+8.86)
+Pringles 48.98 71.45 (+22.47) 73.10 (+24.12) 65.76 (+16.78) 80.69 (+31.71)
+Tincase 55.80 74.82 (+19.02) 79.64 (+23.84) 72.60 (+16.80) 83.96 (+28.16)
+Sandwich 78.36 79.00 (+0.64) 80.41 (+2.05) 78.95 (+0.59) 81.15 (+2.79)
+Mouse 49.85 63.24 (+13.39) 67.61 (+17.76) 58.07 (+8.22) 72.90 (+23.05)
+Overall 62.68 75.07 (+12.39) 77.01 (+14.33) 71.61 (+8.93) 81.59 (+18.91)
+B25
+Spray 76.00 85.90 (+9.90) 77.88 (+1.88) 82.29 (+6.29) 89.49 (+13.49)
+Pringles 64.29 73.93 (+9.64) 73.52 (+9.23) 66.95 (+2.66) 81.17 (+16.88)
+Tincase 74.16 78.98 (+4.82) 81.73 (+7.57) 78.22 (+4.06) 84.71 (+10.55)
+Sandwich 78.10 78.74 (+0.64) 80.15 (+2.05) 78.44 (+0.34) 80.87 (+2.77)
+Mouse 54.54 62.59 (+8.05) 67.12 (+12.58) 60.98 (+6.44) 72.49 (+17.95)
+Overall 69.42 76.03 (+6.61) 76.08 (+6.66) 73.38 (+3.96) 81.75 (+12.33)
+B50
+Spray 76.63 84.29 (+7.66) 82.10 (+5.47) 83.59 (+6.96) 89.56 (+12.93)
+Pringles 67.45 74.74 (+7.29) 76.86 (+9.41) 71.26 (+3.81) 81.43 (+13.98)
+Tincase 78.33 81.24 (+2.91) 82.13 (+3.80) 78.42 (+0.09) 85.00 (+6.67)
+Sandwich 78.38 79.18 (+0.80) 79.90 (+1.52) 78.59 (+0.21) 81.54 (+3.16)
+Mouse 54.98 62.76 (+7.78) 67.39 (+12.41) 60.07 (+5.09) 72.10 (+17.12)
+Overall 71.16 76.44 (+5.28) 77.68 (+6.52) 74.39 (+3.23) 81.93 (+10.77)
+B75
+Spray 80.83 85.88 (+5.05) 83.95 (+3.12) 84.37 (+3.54) 87.20 (+6.37)
+Pringles 67.48 75.38 (+7.90) 76.83 (+9.35) 71.02 (+3.54) 80.31 (+12.83)
+Tincase 75.67 80.73 (+5.06) 82.93 (+7.26) 77.87 (+2.20) 84.44 (+8.77)
+Sandwich 78.38 78.77 (+0.39) 80.08 (+1.70) 78.97 (+0.59) 80.82 (+2.44)
+Mouse 53.93 59.54 (+5.61) 67.66 (+13.73) 59.98 (+6.05) 71.51 (+17.58)
+Overall 71.26 76.06 (+4.80) 78.29 (+7.03) 74.44 (+3.18) 80.86 (+9.60)
+B100
+Spray 83.59 85.20 (+1.61) 85.59 (+2.00) 84.02 (+0.43) 87.22 (+3.63)
+Pringles 69.21 74.31 (+5.10) 76.05 (+6.84) 71.07 (+1.86) 80.55 (+11.34)
+Tincase 77.36 79.33 (+1.97) 79.16 (+1.80) 77.49 (+0.13) 81.20 (+3.84)
+Sandwich 78.74 79.28 (+0.54) 80.21 (+1.47) 79.05 (+0.31) 80.97 (+2.23)
+Mouse 55.80 61.61 (+5.81) 68.22 (+12.42) 60.49 (+4.69) 72.66 (+16.86)
+Overall 72.94 75.95 (+3.01) 77.84 (+4.90) 74.42 (+1.48) 80.52 (+7.58)
+Table 13: AUC@[0:0.1] performance under diverse environments of various sensor control methods for FoundationPose.
+Environment Object
+Baseline
+RGB: Auto
+Depth: Default
+Depth-Only
+RGB: Auto
+Depth: Oracle
+RGB-Only
+RGB: Oracle
+Depth: Default
+Oracle-Fixed
+Best Fixed Param.
+Oracle-Dynamic
+RGB: Oracle
+Depth: Oracle
+B5
+Spray 82.80 84.68 (+1.88) 85.34 (+2.54) 83.44 (+0.64) 87.17 (+4.37)
+Pringles 25.07 39.45 (+14.38) 45.26 (+20.19) 32.98 (+7.91) 64.10 (+39.03)
+Tincase 30.87 49.44 (+18.57) 51.80 (+20.93) 41.60 (+10.73) 65.73 (+34.86)
+Sandwich 81.62 84.41 (+2.79) 83.38 (+1.76) 82.77 (+1.15) 86.28 (+4.66)
+Mouse 67.88 72.80 (+4.92) 75.10 (+7.22) 72.02 (+4.14) 79.54 (+11.66)
+Overall 57.65 66.16 (+8.51) 68.18 (+10.53) 62.56 (+4.91) 76.56 (+18.91)
+B25
+Spray 84.85 86.37 (+1.52) 87.17 (+2.32) 84.90 (+0.05) 88.46 (+3.61)
+Pringles 40.24 51.57 (+11.33) 51.36 (+11.12) 44.67 (+4.43) 71.07 (+30.83)
+Tincase 41.24 63.33 (+22.09) 74.76 (+33.52) 47.56 (+6.32) 84.11 (+42.87)
+Sandwich 83.36 86.72 (+3.36) 84.90 (+1.54) 84.10 (+0.74) 87.97 (+4.61)
+Mouse 71.39 75.37 (+3.98) 75.98 (+4.59) 72.66 (+1.27) 80.51 (+9.12)
+Overall 64.22 72.67 (+8.45) 74.83 (+10.61) 66.78 (+2.56) 82.43 (+18.21)
+B50
+Spray 85.49 86.85 (+1.36) 88.22 (+2.73) 85.32 (-0.17) 89.59 (+4.10)
+Pringles 35.67 55.31 (+19.64) 50.74 (+15.07) 45.50 (+9.83) 79.00 (+43.33)
+Tincase 47.24 63.93 (+16.69) 71.60 (+24.36) 54.20 (+6.96) 82.71 (+35.47)
+Sandwich 83.62 87.13 (+3.51) 85.46 (+1.84) 85.00 (+1.38) 88.56 (+4.94)
+Mouse 70.71 74.39 (+3.68) 75.78 (+5.07) 72.51 (+1.80) 79.88 (+9.17)
+Overall 64.54 73.52 (+8.98) 74.36 (+9.82) 68.51 (+3.97) 83.95 (+19.41)
+B75
+Spray 85.93 87.20 (+1.27) 88.73 (+2.80) 86.12 (+0.19) 89.83 (+3.90)
+Pringles 35.19 49.95 (+14.76) 60.00 (+24.81) 54.76 (+19.57) 81.88 (+46.69)
+Tincase 39.51 59.67 (+20.16) 78.89 (+39.38) 52.82 (+13.31) 86.93 (+47.42)
+Sandwich 83.56 86.82 (+3.26) 85.82 (+2.26) 84.92 (+1.36) 88.92 (+5.36)
+Mouse 70.98 74.78 (+3.80) 75.59 (+4.61) 72.88 (+1.90) 80.10 (+9.12)
+Overall 63.03 71.68 (+8.65) 77.81 (+14.78) 70.30 (+7.27) 85.53 (+22.50)
+B100
+Spray 84.07 87.51 (+3.44) 88.54 (+4.47) 86.00 (+1.93) 90.00 (+5.93)
+Pringles 35.14 57.31 (+22.17) 46.55 (+11.41) 48.07 (+12.93) 77.40 (+42.26)
+Tincase 38.89 55.69 (+16.80) 73.11 (+34.22) 57.33 (+18.44) 85.27 (+46.38)
+Sandwich 83.69 86.92 (+3.23) 86.03 (+2.34) 85.21 (+1.52) 88.82 (+5.13)
+Mouse 70.17 74.37 (+4.20) 76.88 (+6.71) 72.46 (+2.29) 79.85 (+9.68)
+Overall 62.39 72.36 (+9.97) 74.22 (+11.83) 69.81 (+7.42) 84.27 (+21.88)
+Appendix G The Use of Large Language Models (LLMs)
+In this work, editing, data processing, and visualization were polished with the assistance of ChatGPT (OpenAI, 2025) and Gemini (Google, 2025).